### v2

# Improving the Quality of Web-mined Parallel Corpora of Low-Resource Languages using Debiasing Heuristics

Abstract
Parallel Data Curation (PDC) techniques aim to filter out noisy parallel sentences from web-mined corpora. Ranking sentence pairs using similarity scores on sentence embeddings derived from Pre-trained Multilingual Language Models (multiPLMs) is the most common PDC technique. However, previous research has shown that the choice of the multiPLM significantly impacts the quality of the filtered parallel corpus, and the NMT models trained using such data show a disparity across multiPLMs. This paper shows that this disparity is due to different multiPLMs being biased towards certain types of sentence pairs, which are treated as noise from an NMT point of view. We show that such noisy parallel sentences can be removed to a certain extent by employing a series of heuristics. The curated corpus leads to producing better results by NMT models while minimizing the disparity across multiPLMs. We publicly release the source code and the curated datasets [^1] Source codes and Datasets will be released upon acceptance.
Improving the Quality of Web-mined Parallel Corpora of Low-Resource Languages using Debiasing Heuristics
Aloka Fernando1, Nisansa de Silva1, Menan Velayuthan1, Charitha Rathnayake1 Surangika Ranathunga 2 {alokaf,NisansaDdS,velayuthan.22,charitha.18}@cse.mrt.ac.lk s.ranathunga@massey.ac.nz 1Dept. of Computer Science and Engineering, University of Moratuwa,10400, Sri Lanka 2Massey University, Palmerston North, New Zealand, 4443 Correspondence: alokaf@cse.mrt.ac.lk
1 Introduction
Parallel data mined from the web at scale is often considered an alternative to human-created data in training Neural Machine Translation (NMT) models Costa-jussà et al. (2022); Bañón et al. (2020). CCAligned El-Kishky et al. (2020), CCMatrix Schwenk et al. (2021) and ParaCrawl Bañón et al. (2020) are examples of such web-mined corpora, which cover Low-Resource Languages (LRLs) as well. However, quality audits of these corpora, as shown in Table 1, reveal substantial noise, which negatively impacts the NMT models Khayrallah and Koehn (2018).
Noise Category A B C D E
Perfect translations - - - Y Y
Near perfect translation - - - - Y
Correct translation - Low quality - Y - Y Y
Over/Under translation - Y Y Y -
Misordered words Y Y Y Y -
Spelling permutations - Y - Y -
Untranslated Sentences Y Y Y - Y
Short Sentences Y - Y - Y
Mismatch numbers Y - - -
Machine Translated Sentences - - Y Y -
Misaligned sentences Y Y Y Y Y
Wrong Language Y Y Y Y Y
Not a Language Y Y Y Y Y
Table 1: Noise types in parallel corpora, as identified by Khayrallah and Koehn (2018) (A), Bane et al. (2022) (B), Herold et al. (2022) (C), Kreutzer et al. (2022) (D) and Ranathunga et al. (2024) (E).
Parallel Data Curation (PDC) aims at improving the quality of such web-mined parallel corpora by filtering high-quality parallel sentence pairs. The importance of PDC for LRLs has been emphasised with the introduction of WMT shared tasks Sloto et al. (2023); Koehn et al. (2020, 2019). Initiated by the work of  Chaudhary et al. (2019), recent PDC techniques follow a scoring and ranking mechanism using embeddings obtained from a Multilingual Language Model (multiPLM). Then, a subset of the top-ranked sentence pairs is selected to train the NMT model. However, the quality of these top-ranked pairs depends on the selected multiPLM (Ranathunga et al., 2024; Moon et al., 2023).
To investigate this further, we obtain embeddings from three multiPLMs: LASER3 Heffernan et al. (2022), XLM-R Conneau et al. (2020), and LaBSE Feng et al. (2022), calculate the semantic similarity between each parallel sentence pair in the CCMatrix Schwenk et al. (2021) and CCAligned El-Kishky et al. (2020) datasets and rank them in descending order. Then we train NMT models (Section 4.4) using the top-ranked 100k sentence pairs from each corpus. Experiments are carried out for English-Sinhala (En-Si), English-Tamil (En-Ta) and Sinhala(Si-Ta). As shown in Figure 1, there is a significant disparity among the results, mainly for En-Si and En-Ta language pairs.
Figure 1: Baseline NMT scores using top-ranked sentence pairs from CCMatrix and CCAligned, using embeddings obtained from LASER3, XLM-R, and LaBSE.
A manual inspection of the En-Si top-ranked 1000 sentences reveal that different multiPLMs prioritise different sentence characteristics when ranking parallel sentences. For example, sentence pairs ranked top by XLM-R are mostly short and contain overlapping text such as numbers, acronyms, and URLs. LaBSE also favours sentences with numbers and date overlaps, while LASER3 prioratizes relatively better full length sentence-pairs. The impact of the nature of the top-ranked sentences is evident in Figure 1, where results related to LASER3 outperform its two counterparts by a significant margin, in most of the experiments.
In order to carry out a more quantitative evaluation, we randomly selected 100 sentence pairs from the top 1000 sentences from the aforementioned ranked corpora and carried out a human evaluation. Ranathunga et al. (2024)’s is the most comprehensive error taxonomy available for this task. However, to better capture the noise identified during the manual observation, we extended this error taxonomy (Section 3). Human evaluation results in Table 2 show that compared to LASER3, the error category percentage (E) is excessively high for XLM-R and LaBSE. The most contributing types of errors are untranslated text (UN), short sentences (CS) and sentences with number/acronym/URL overlaps (CCN). Examples for these are shown in Table 10 in Appendix F.
$C$ CN CB     C     CS CCN UN X WL NL     E     $C$ CN CB     C     CS CCN UN X WL NL     E     $C$ CN CB     C     CS CCN UN X WL NL     E
Sinhala - Tamil English - Sinhala English - Tamil
CCMatrix
LASER3-Before 8.00% 27.00% 2.00%     37.00%     14.00% 14.00% 34.00% 1.00% 0.00% 0.00%     63.00%     17.00% 7.00% 4.00%     28.00%     7.00% 10.00% 55.00% 0.00% 0.00% 0.00%     72.00%     0.00% 3.00% 2.00%     5.00%     0.00% 0.00% 95.00% 0.00% 0.00% 0.00%     95.00%
LASER3-After 16.00% 56.00% 1.00%     73.00%     13.00% 4.00% 10.00% 0.00% 0.00% 0.00%     27.00%     39.00% 39.00% 7.00%     85.00%     0.00% 7.00% 8.00% 0.00% 0.00% 0.00%     15.00%     6.00% 61.00% 20.00%     87.00%     0.00% 3.00% 10.00% 0.00% 0.00% 0.00%     13.00%
XLM-R-Before 1.00% 10.00% 0.00%     11.00%     40.00% 19.00% 29.00% 0.00% 1.00% 0.00%     89.00%     1.00% 0.00% 0.00%     1.00%     13.00% 4.00% 80.00% 2.00% 0.00% 0.00%     99.00%     0.00% 0.00% 2.00%     2.00%     3.00% 5.00% 90.00% 0.00% 0.00% 0.00%     98.00%
XLM-R - After 0.00% 20.00% 2.00%     22.00%     13.00% 29.00% 35.00% 0.00% 1.00% 0.00%     78.00%     3.00% 8.00% 26.00%     37.00%     0.00% 2.00% 53.00% 8.00% 0.00% 0.00%     63.00%     0.00% 39.00% 31.00%     70.00%     1.00% 3.00% 21.00% 4.00% 0.00% 1.00%     30.00%
LaBSE - Before 4.00% 6.00% 0.00%     10.00%     74.00% 7.00% 9.00% 0.00% 0.00% 0.00%     90.00%     13.00% 2.00% 0.00%     15.00%     63.00% 14.00% 8.00% 0.00% 0.00% 0.00%     85.00%     0.00% 9.00% 2.00%     11.00%     34.00% 7.00% 48.00% 0.00% 0.00% 0.00%     89.00%
LaBSE - After 29.00% 33.00% 0.00%     62.00%     2.00% 32.00% 4.00% 0.00% 0.00% 0.00%     38.00%     87.00% 7.00% 3.00%     97.00%     0.00% 1.00% 2.00% 0.00% 0.00% 0.00%     3.00%     36.00% 53.00% 4.00%     93.00%     1.00% 3.00% 2.00% 1.00% 0.00% 0.00%     7.00%
CCAligned
LASER3-Before 3.00% 24.00% 3.00%     30.00%     34.00% 19.00% 17.00% 0.00% 0.00% 0.00%     70.00%     2.00% 22.00% 8.00%     32.00%     13.00% 30.00% 23.00% 2.00% 0.00% 0.00%     68.00%     2.00% 23.00% 18.00%     43.00%     13.00% 27.00% 17.00% 0.00% 0.00% 0.00%     57.00%
LASER3-After 5.00% 79.00% 2.00%     86.00%     0.00% 9.00% 4.00% 1.00% 0.00% 0.00%     14.00%     13.00% 58.00% 14.00%     85.00%     0.00% 0.00% 13.00% 2.00% 0.00% 0.00%     15.00%     3.00% 67.00% 10.00%     80.00%     0.00% 8.00% 12.00% 0.00% 0.00% 0.00%     20.00%
XLM-R-Before 0.00% 0.00% 2.00%     2.00%     48.00% 49.00% 0.00% 0.00% 0.00% 1.00%     98.00%     2.00% 0.00% 0.00%     2.00%     72.00% 20.00% 6.00% 0.00% 0.00% 0.00%     98.00%     0.00% 8.00% 4.00%     12.00%     42.00% 16.00% 15.00% 8.00% 0.00% 7.00%     88.00%
XLM-R - After 20.00% 33.00% 4.00%     57.00%     1.00% 22.00% 19.00% 0.00% 1.00% 0.00%     43.00%     18.00% 18.00% 20.00%     56.00%     0.00% 6.00% 34.00% 4.00% 0.00% 0.00%     44.00%     6.00% 46.00% 30.00%     82.00%     0.00% 9.00% 9.00% 0.00% 0.00% 0.00%     18.00%
LaBSE - Before 0.00% 1.00% 0.00%     1.00%     69.00% 26.00% 3.00% 0.00% 0.00% 1.00%     99.00%     0.00% 1.00% 0.00%     1.00%     97.00% 2.00% 0.00% 0.00% 0.00% 0.00%     99.00%     0.00% 1.00% 0.00%     1.00%     97.00% 0.00% 0.00% 0.00% 0.00% 2.00%     99.00%
LaBSE - After 15.00% 34.00% 0.00%     49.00%     2.00% 43.00% 6.00% 0.00% 0.00% 0.00%     51.00%     45.00% 27.00% 3.00%     75.00%     1.00% 19.00% 5.00% 0.00% 0.00% 0.00%     25.00%     19.00% 45.00% 3.00%     67.00%     0.00% 22.00% 11.00% 0.00% 0.00% 0.00%     33.00%
Table 2: Human evaluation results for CCMatrix and CCAligned for En-Si, En-Ta and Si-Ta. Results are reported for LASER3, XLM-R, and LaBSE before and after applying heuristics. (C) - overall correct percentage considering $C$ (perfect translation), CN (near perfect) and CB (boilerplate). (E) - overall error percentage considering CCN (correct with overlaps), CS (correct but short sentence), X (wrong translation), UN (untranslated), WL (wrong language), NL (not a language).
We hypothesis that at least some of these noisy sentences can be filtered using rule-based heuristics. Although applying heuristics is a common approach to improving the quality of parallel corpora Sloto et al. (2023); Steingr´ımsson et al. (2023), the use of heuristics has not been consistent. They have limited their analysis to a single commonly used heuristic, or have varied the criteria or thresholds of the heuristics arbitrarily from one study to another. Further, their work does not explore the potential of combining multiple heuristics.
In this research, we incorporate heuristics proposed in previous studies, along with a new heuristic of our own, and conduct an empirical study to investigate whether a more refined selection of top-ranked parallel sentences can be identified by systematically combining these heuristics.
Our key contributions are as follows:
  • We extend  Ranathunga et al. (2024)’s error taxonomy with a new error category to capture additional type of noise, which refers to sentence pairs with high-overlapping instances of numbers, acronyms, URLs.
  • We conduct a systematic study to analyse the impact of rule-based heuristics in filtering noisy sentences from the web-mined corpora and identify an optimal combination of heuristics that works across the corpora and languages considered in the study.
  • We empirically show that applying heuristics before ranking sentences based on embeddings derived from multiPLMs results in higher NMT scores, and reduces the disparity across multiPLMs.
  • We conduct a human evaluation to assess the impact of noise filtering across three multiPLMs.
2 Related Work
2.1 MultiPLMs for PDC
While employing a multiPLM for PDC is common, past research experimented with only one multiPLM at a time. For example, in the WMT2023 shared task, LASER2 was utilized to set the task baseline, whereas Steingr´ımsson (2023) and Gala et al. (2023) only used LaBSE. Therefore, the disparities across multiPLMs and biases specific to each multiPLM have not come into light.
On the other hand, studies conducted by Ranathunga et al. (2024) and Moon et al. (2023) reveal that using different multiPLMs for scoring and ranking parallel corpora, and training NMT models with the top-ranked corpora, results in a disparity. Moon et al. (2023) observe that this is due to biases in multiPLMs, which tend to rank noisy parallel sentences highly. However, there has been no study to identify these biases.
2.2 Identifying Noise in Web-mined Corpora
Recent research used categorical labels to annotate translation pairs, aiming at quantifying the noise types in web-mined corpora. Kreutzer et al. (2022) used their taxonomy to conduct manual audits on random samples from three web-mined datasets and reported substantial noise, specifically for LRLs. Ranathunga et al. (2024)’s taxonomy (See Table 9 in Appendix F) is an extension of Kreutzer et al. (2022)’s taxonomy. They first ranked the datasets based on embeddings from a multiPLM, and then selected random samples from top and bottom portions. Their human evaluation based on the taxonomy reported that the quality of the parallel sentences varies heavily depending on the selected portion.
2.3 Heuristic-based PDC
Commonly used rule-based heuristics can be categorised into four groups as described below:
De-duplication based (Dedup): Removing identical duplicates in the monolingual sides is a common practice Costa-jussà et al. (2022), while De-duplicating after removing non-alpha characters and punctuations Bala Das et al. (2023) could be found as its variants.
Length-based (sLength) : Gala et al. (2023) and Aulamo et al. (2023) have identified the removal of short sentences as a potential heuristic. Short sentences hinder NMT models in two ways Koehn and Knowles (2017): by providing insufficient syntactic and semantic information, or resulting in an overfitting situation.
LID-based (LID): Language Identification is used to remove fully/partially untranslated text and content in the wrong language Steingr´ımsson et al. (2023); Gala et al. (2023); Zhang et al. (2020).
Ratio-based : Ratio-based heuristics identify and remove sentences that show significant structural imbalances between the source and target sentences. It is based on the assumption that, well-aligned sentence pairs tend to maintain consistent ratios in terms of character count, word count, or token distribution. We observe three common types of ratio-based heuristics: (1) source-to-target sentence length ratio (STRatio(Rossenbach et al., 2018; Gale and Church, 1993), (2) alpha-only words to sentence words ratio (sentWRatio(Aulamo et al., 2020) and (3) alpha-only character ratio with respect to the sentence characters (sentCRatioHangya and Fraser (2018).
However, the impact of these heuristics in isolation and as a combination had not been evaluated systematically.
Ranathunga et al. (2024)
Improved
Taxonomy
Revision
CCN Perfect/near perfect translation where more than 30% of the overlapping content is numbers/acronyms/URLs/etc
Short Sentences (Max 3 words) CS Less than 5 words on either side
Wrong Language WL Specifically set a threshold as 30%
Table 3: A comparison of the improved taxonomy against Ranathunga et al. (2024)’s (only showing the changes).
3 Methodology
Improved Taxonomy for Noise Although translation pairs can have overlapping URLs, acronyms, etc, excessive inclusions of such content in a sentence (e.g. consider the sentence ‘Contact: Diane Anderson 076-8268914, info@sandnasbadenscamping.se’ [^2] More examples are in Table 8 in Appendix A) do not provide meaningful content for an NMT system to learn from. However, under  Ranathunga et al. (2024)’s taxonomy, such sentence pairs would likely be categorized as correct translations. Therefore, we define a new noise category CCN to capture such sentence pairs.
Secondly, we consider the upper limit for short sentences as five words [^3] We consider thresholds 3,4 and 5 and empirically select this threshold as the optimal., based on common observations. Finally, we improve the definition of WL (wrong language) to consider a threshold in determining whether a sentence pair should be marked as wrong language. Table 3 shows these changes. The complete list of these noise categories and example parallel sentences is available in Table 9 and Table 10 (respectively) in Appendix A.
Selection of Heuristics
Table 4 shows how the heuristics discussed in Section 2.3 may help in removing different noise categories. Note that de-duplication based heuristic cannot be associated with any noise category, as it does not apply to individual sentence pairs. In addition to de-duplication strategies discussed in Section 2.3, we introduce an n-gram based de-duplication, meaning that sentences would be removed if they overlap in a consecutive n-gram sequence.
Noise Category Short Label Rule-based Heuristic
Not a language NL LID, sentWRatio, sentCRatio
Wrong language WL LID
Untranslated UN LID
Short Sentences CS sLength
Overlapping Numbers/acronyms/URLs/email CCN LID, sentWRatio, sentCRatio
Wrong translation X STRatio (With a length difference)
Boilerplate translation CB STRatio
Table 4: Mapping between the noise category vs the noise mitigating heuristic.
Human Evaluation
We conduct a human evaluation to quantify the noise before and after applying the heuristics. The annotator selection criteria, resources and training provided, the payment details, etc are described in Appendix F.
From the top 1000 samples in each of the ranked corpora using LASER3, XLM-R and LaBSE, we randomly select 100 parallel sentences for each language pair. We ask each translator to select an annotation category for each sentence pair using the error taxonomy discussed in Section 3. Each sentence pair was annotated by two translators to reduce potential bias caused by the individual translators. Inter-annotator agreement was calculated using the Pearson correlation coefficient, which resulted in En-Si 0.88, En-Ta 0.68 and Si-Ta 0.49.
4 Experiments
4.1 Data
We use the language pairs, En-Si, En-Ta, and Si-Ta in our experiments. Sinhala and Tamil are morphologically rich, low-resource and mid-resourced languages Joshi et al. (2020); Ranathunga and De Silva (2022), respectively. Languages were selected considering the availability of human evaluators. We select CCMatrix and CCAligned as the web-mined corpora. Both these corpora include parallel data for the language pairs considered in the search. More details of these datasets and language pairs are in Appendix B.
For the NMT experiments, we use the dev and devtest subsets from the Flores+ dataset [^4] https://github.com/openlanguagedata/flores as validation and evaluation sets, respectively. Dataset statistics are provided in Table 5.
Language-pair CCMatrix CCAligned dev devtest
En-Si 6,270,801 619,711 997 1,012
En-Ta 7,291,119 880,547 997 1,012
Si-Ta 215,966 260,118 997 1,012
Table 5: Corpus statistics.
4.2 Selection of muliPLMs
We select LASER3, XLM-R, and LaBSE for obtaining embeddings for the sentences to determine the semantic similarity. XLM-R, different to others, was trained purely on monolingual data, but has proven to be useful for cross-lingual tasks as well Choi et al. (2021); Conneau et al. (2020). All three models include En, Si, and Ta. Details of the multiPLMs are shown in Appendix C.
4.3 Heuristic-based PDC Experiments
Each heuristic is applied independently to the source (S), target (T), and both sides (ST). In line with the original sentence alignment conducted for CCMatrix and CCAligned, we treat En as the source side. For Si-Ta, Si is considered the source because it is more common for a Si sentence to be translated to Ta Farhath et al. (2018). Finally, for each multiPLM, the retained sentences are ranked in descending order with cosine similarity.
Deduplication: We consider different granularities of de-duplication. ie. identical de-duplication (dedup), de-duplicate by removing numbers only (nums) and removing both numbers and punctuations (punctsNums). Subsequently, we de-duplicate considering different n-gram spans, i.e. 4-grams, 5-grams, 6-grams and 7-grams.
Length-based: We filter short sentences less than five words [^5] We consider sentence length thresholds 3,4 and 5 and empirically decide this value. Results in Table 11 in Appendix D. While some research has suggested removing extremely long sentences (Minh-Cong et al., 2023; Gala et al., 2023), we found that the percentage of longer sentences is lower and that removing them has a negligible effect. Thus, this result is not reported.
LID-based: We use a public LID model [^6] https://github.com/facebookresearch/fairseq/tree/nllb to predict the language of each sentence. The predicted label is then used as a standalone heuristic (LID) and in combination with its associated prediction probability (LIDThresh), with threshold of 0.7777Thresholds below 0.7 reduce NMT results.
Ratio-based: For STRatio, 0.79-1.39, 0.87-1.62 and 0.85-1.57 were selected as thresholds for En-Si, En-Ta and Si-Ta, respectively. These were determined by calculating the mean and the standard deviation obtained for the validation set in a human-crafted En-Si-Ta trilingual dataset (Fernando et al., 2020). Following observations of Hangya and Fraser (2018), 0.6 was selected as the threshold for sentWRatio and  sentCRatio.
4.4 NMT Experiments
First, a Sentencepiece [^8] https://github.com/google/sentencepiece tokenizer with a vocabulary size of 25000 is trained. Then we use the fairseq toolkit  (Ott et al., 2019) to model and train the transformer-based Seq-to-Seq NMT model until convergence. Hyperparameters used in the NMT experiments are shown in Table 12 in Appendix E. The baseline NMT models are trained on the top 100,000 sentence pairs from the ranked corpus. We use ChrF++ Popović (2017) to report NMT results.
5 Results and Analysis
We report the results obtained for the NMT models trained in the forward direction. The results of the experiments are in Table 6. Heuristic-wise best result is summarized in Table 14 in Appendix G. As evident from these tables, as well as from Figure 1, NMT results across different multiPLMs show a great disparity in the baseline NMT scores for En-Si and En-Ta language pairs. In the following sub-sections, we discuss how the use of heuristics is useful in mitigating this disparity and improving overall NMT results.
5.1 Impact of Heuristics on NMT Results
Heuristic Applicable Side Sinhala-Tamil English-Sinhala English-Tamil
CCMatrix CCAligned CCMatrix CCAligned CCMatrix CCAligned
LASER3 XLM-R LaBSE LASER3 XLM-R LaBSE LASER3 XLM-R LaBSE LASER3 XLM-R LaBSE LASER3 XLM-R LaBSE LASER3 XLM-R LaBSE
Baseline 31.08 30.99 31.63 35.36 35.97 35.79 30.76 5.55 14.49 32.33 19.39 27.57 19.02 5.86 14.20 40.13 17.40 26.00
$D$ S 32.05 31.50 32.07 36.40 36.01 34.98 29.72 6.35 14.69 33.26 21.04 28.22 19.67 4.93 14.96 40.87 19.47 26.26
T 31.39 31.44 31.73 36.26 35.86 35.96 33.81 12.59 25.97 33.66 21.41 28.32 19.48 6.87 17.96 40.13 17.90 27.79
ST 32.26 31.10 32.25 36.41 36.08 35.32 34.01 13.80 26.18 33.47 22.22 29.49 20.32 6.45 17.53 40.56 19.83 30.01
DD-4gram S 30.37 30.65 30.53 35.74 35.24 34.55 28.69 8.56 13.05 31.56 23.53 28.25 19.72 7.06 19.56 35.54 25.64 26.49
T 31.00 29.90 29.39 36.05 35.98 35.44 31.79 13.60 23.66 32.86 24.95 29.05 19.82 7.08 20.23 39.83 27.44 31.18
ST 30.86 31.13 30.80 35.28 35.36 34.64 28.72 15.17 20.45 28.15 15.45 21.37 18.15 7.00 21.37 35.02 25.70 27.41
DD-5gram S 30.89 30.90 31.25 35.64 35.81 35.87 28.73 7.14 13.51 33.44 23.98 28.79 18.06 4.70 17.16 40.39 24.07 29.07
T 31.24 31.55 32.10 36.26 35.87 35.23 33.98 14.01 26.23 34.10 22.27 31.10 20.15 6.75 18.78 41.12 24.05 30.26
ST 30.78 31.53 31.35 35.64 35.94 35.44 31.95 13.87 23.07 31.60 17.10 23.52 19.61 6.25 20.12 21.77 25.22 29.36
DD-6gram S 31.89 30.82 31.76 36.31 36.11 35.88 31.10 7.62 13.41 33.53 21.47 28.51 20.32 5.47 15.59 40.48 21.75 27.64
T 32.51 30.41 32.29 36.35 36.23 36.01 34.21 13.98 24.91 34.24 23.63 30.23 21.75 6.69 20.32 40.44 20.31 30.48
ST 31.89 30.82 31.76 35.84 35.95 35.54 33.63 14.96 24.72 33.29 15.54 25.55 20.38 7.18 20.19 41.73 24.89 31.06
DD-7gram S 31.48 31.27 32.03 36.26 35.67 35.50 30.93 5.91 15.94 33.27 19.90 29.58 21.54 5.71 16.49 40.63 20.01 28.91
T 31.56 31.06 30.85 36.44 36.10 35.16 34.27 13.72 25.58 32.97 22.14 28.22 20.91 7.37 21.96 40.49 19.18 28.69
ST 31.48 31.27 32.03 35.74 35.90 34.82 33.93 14.95 24.95 33.63 14.58 24.96 17.56 5.98 20.71 40.94 22.16 29.40
DD+N S 31.51 31.37 31.99 36.61 36.66 35.99 30.54 5.92 15.12 34.77 28.07 31.81 17.00 5.60 13.41 41.40 28.65 35.22
T 31.17 30.51 32.09 36.30 36.45 36.32 33.83 14.44 25.86 34.47 27.27 31.90 17.54 6.09 19.01 41.36 28.40 35.12
ST 31.71 31.22 31.66 36.49 36.37 36.10 33.83 14.15 26.12 34.24 28.45 31.64 19.19 5.15 18.92 41.46 30.49 35.42
DD+PN S 31.90 31.47 31.02 36.50 36.00 36.12 30.55 6.28 16.67 34.72 27.25 31.89 18.15 5.79 15.66 41.78 30.55 35.78
T 31.90 32.05 30.89 36.63 36.47 36.86 33.89 14.81 26.31 35.06 27.69 32.01 21.57 8.24 20.41 41.64 29.35 35.32
ST 32.05 31.31 32.53 35.96 36.71 36.23 33.37 14.15 26.08 34.08 27.80 32.59 20.99 5.82 18.83 41.80 30.69 35.91
DD+PN+4gram ST+T NA NA NA 30.64 29.48 30.19 NA 41.82 35.90 37.08
DD+PN+5gram ST + T 32.98 32.73 32.60 36.24 36.21 36.35 34.50 16.09 25.78 33.81 30.33 32.74 NA NA
DD+PN+6gram ST + T 30.41 31.38 31.42 36.73 36.62 36.37 NA 35.24 28.21 31.26 19.49 6.67 20.60 41.90 35.97 35.94
DD+PN+7gram T +T NA NA NA NA 19.57 7.55 20.89
SL S 31.41 31.52 32.30 36.42 36.37 36.52 32.49 6.58 20.70 33.86 26.53 32.97 17.50 5.11 18.74 41.40 27.60 36.77
T 31.38 30.56 31.97 36.30 36.71 36.58 31.88 7.83 28.51 34.88 29.42 33.14 18.52 6.33 21.73 41.54 30.16 37.61
ST 31.21 31.32 31.37 36.47 35.99 36.60 32.82 8.24 29.96 34.83 29.55 33.50 19.45 5.33 20.79 41.14 32.67 38.08
LID S 31.48 31.36 31.78 36.05 36.03 35.64 31.00 6.23 14.69 34.39 27.33 31.73 18.44 6.93 13.43 41.80 31.41 33.95
T 30.78 31.14 31.53 35.68 36.07 35.85 32.48 12.22 16.04 33.70 24.38 30.48 29.59 14.70 24.24 41.51 24.24 30.69
ST 31.43 30.66 31.40 36.17 36.12 35.18 31.99 13.32 16.20 34.11 28.87 32.26 29.59 13.54 23.45 41.42 32.33 36.13
LT S 30.05 31.25 31.06 35.60 35.25 34.29 30.32 7.12 15.26 35.73 30.86 32.69 18.98 6.02 13.06 41.60 35.25 36.29
T 31.28 30.40 30.68 35.03 30.01 32.01 32.82 12.94 15.81 35.22 27.46 30.40 29.59 15.24 24.51 41.03 30.01 34.01
ST 30.33 30.46 30.71 36.73 36.73 36.80 32.84 14.08 13.71 35.11 32.97 32.88 28.93 15.16 25.33 42.63 38.01 37.40
STRatio - 31.74 22.80 31.34 36.39 35.74 35.30 31.09 5.20 15.40 33.47 24.05 30.21 20.52 5.40 18.29 40.91 22.71 28.61
sentWRatio S 30.65 30.62 32.03 36.17 35.77 35.54 31.50 7.40 10.86 34.15 25.97 31.35 19.42 5.79 13.93 42.05 29.70 35.53
T 30.71 31.59 31.34 36.24 36.17 36.46 30.99 6.39 15.13 33.51 26.93 30.47 18.61 5.65 11.08 41.87 30.06 35.54
ST 31.93 31.56 30.98 36.44 36.72 36.01 30.64 7.00 15.50 33.85 28.73 31.17 18.99 4.82 14.08 41.05 30.88 35.77
sentCRatio S 31.67 31.24 31.14 35.94 36.18 35.86 30.15 7.05 14.46 34.06 21.52 30.10 17.47 6.22 13.83 40.68 22.48 29.37
T 30.98 31.21 31.93 36.36 35.43 35.85 30.65 5.83 15.28 33.64 23.14 29.05 19.90 6.78 12.51 40.78 19.63 29.42
ST 32.28 31.90 32.04 36.33 35.60 36.11 30.85 6.45 14.64 33.60 23.84 29.70 19.54 6.45 10.79 41.76 21.82 30.82
Combined Heuristics
DD+PN+ngram (SiTa-CCMatrix n=5, SiTa-CCAligned n= 7 EnSi-CCMatrix/CCAligned n=5, EnTa-CCMatrix n=7, EnTa-CCAligned n=6)
+sLength T + ST 30.17 29.02 29.99 36.32 36.81 36.61 35.03 21.70 26.32 35.68 33.49 34.43 30.29 19.44 29.85 42.84 39.36 40.16
+LT T + ST 31.49 30.13 30.68 36.58 36.37 37.02 35.42 19.58 32.43 34.77 32.58 34.72 20.53 7.52 23.35 42.68 38.45 39.60
+sentWRatio T+S 31.37 30.55 30.92 36.83 36.75 36.30 33.99 15.76 24.92 33.97 31.40 32.72 21.67 8.23 24.58 42.11 37.47 38.07
+SL+LT T + ST 29.28 30.85 29.96 36.47 36.81 36.88 35.70 23.92 32.77 34.97 34.92 35.60 30.65 20.86 31.49 42.85 41.17 41.31
+SL+sentWRatio T + ST + ST 31.45 32.65 31.17 36.60 36.85 36.32 35.71 18.93 32.53 35.45 33.42 33.82 22.46 9.11 23.82 41.97 40.07 40.06
+SL+LT+sentWRatio T+ST+ST+S 29.81 29.53 29.73 36.83 36.66 37.03 36.10 23.84 33.94 36.15 34.50 35.67 NA 43.47 41.74 41.06
+SL+LT+sentWRatio>0.8 T+ST+ST+ST 28.70 28.39 28.34 36.20 36.60 35.89 35.66 24.18 33.19 36.26 35.66 35.42 NA 42.08 40.56 42.02
+SL+LT+sentCRatio T+ST+ST+ST 32.64 31.30 32.28 NA NA NA NA NA
+SL+LT+STRatio T+ST+ST+STR NA NA NA NA 30.67 23.36 31.80 NA
Table 6: NMT results obtained after applying heuristics in isolation and in combination in the ablation study. The values in bold indicate the highest NMT score obtained for a given heuristic class or from the heuristic combination. The values underlined are the highest among the individual heuristics. Highlighted in green are the overall best values. Here DD+PN is Deduplication+punctNums, SL is sLength and LT is LIDThresh. NA would be when the particular experiment is not applicable for that language pair or the dataset.
5.1.1 Impact of De-duplication based PDC
We observe that dedup, irrespective of the heuristic-applied side (S/T/ST), outperforms the baseline 89% of the experiments. Overall, de-duplication considering both source and target seems to be the most effective - it outperforms the baseline in 94% of the experiments. In comparison, dedup target only and source only outperform the baseline in 89% and 83% of the experiments, respectively.
We apply our newly introduced ngram-based de-duplication dedup+ngram, on top of  dedup. We find that for each result column, there is a dedup+ngram result that outperforms the best results obtained with the corresponding dedup result. To observe the impact of n value on the NMT result, we plot Figure 2 showing the percentage of ngram experiments exceeding the highest dedup, with respect to the multiPLM (language-wise result is in Figure 4 of the Appendix G). We observe a consistent pattern - n=5 or 6 perform the best in a majority of cases. We believe that 4-gram results in an overly aggressive de-duplication. However, the exact ngram depends on the corpus characteristics.
Figure 2: Percentage of dedup+ngram experiments exceeding the best result of dedup for each multiPLM
We observe that dedup+punctNums outperforms  dedup+nums and dedup in 78% and 89% of the experiments (respectively), proving that (dedup+punctNums) to be more impactful. Finally, we analyse the impact of dedup+puntsNums+ngram999 puntsNums and ngram has been applied on top of dedup.. Compared to other dedup combinations101010dedup, dedup+ngram, nums, and dedup+punctNums, dedup+punctNums+ngram produce the best result across 67% of the experiments.
5.1.2 Impact of Length-Based PDC
sLength surpasses the baseline 89% of the experiments. Therefore, we can conclude that sLength is favourable as a heuristic. When analyzing the side on which the heuristic is applied, we observe that applying it to both ST is the most effective (similar to dedup), followed by T and then S (56%, 28%, and 17% of the experiments, respectively).
Recall that our manual inspection noticed that XLM-R and LaBSE tend to prioritize shorter sentences. This observation is affirmed by the sLength results. For En-Si and En-Ta, this heuristic resulted in substantial gains for many of the XLM-R and LaBSE experiments, while it shows marginal improvements for LASER3.
5.1.3 Impact of LID-Based PDC
With LID-based PDC, we observe substantial improvements for XLM-R and LaBSE, except for Si-Ta, for which the gains are marginal. Gains are reported by LASER3 as well, though not very significant. The highest gain of +20.61 ChrF++ is reported for XLMR for the CCAligned-EnTa corpus, while a gain of +11.40 is reported by LaBSE for the same corpus.
NMT models trained after applying  LID outperform the baseline in 85% of the experiments, while  LIDThresh outperforms LID in 72% of the experiments. Therefore, we conclude that LIDThresh would be the most suitable heuristic. We observe that the gains for LIDThresh with respect to LID are least with Si-Ta with 50% while for En-Si and En-Ta it is 83% each. We assume this is due to a limitation with the LID model, which is not optimized for Si and Ta.
5.1.4 Impact of Ratio-based PDC
We observe that STRatiosentWRatio and sentCRatio produce NMT gains over baseline for 56%, 69% and 80% of experiments, respectively. Among the three ratio-based heuristics, sentWRatio outperforms both STRatio and sentCRatio in 67% of the experiments, making it the most effective for noise reduction. In contrast, STRatio and sentCRatio exceed the performance of the other two heuristics in only 11% and 22% of the cases, respectively. Maximum gains are reported for the EnTa-CCAligned corpus, with +1.92, +13.48, and +9.77 ChrF++ for LASER3, XLM-R, and LaBSE respectively.
5.2 Summary of Heuristic-based PDC
1. Impact of the Individual Heuristics on the NMT Results: The LID-based heuristic emerged as the most impactful in 44% of the experiments. De-duplication and sentence-length heuristics were the most effective in 33% and 17% of the experiments, respectively. The ratio-based heuristic alone did not hold superior results compared to others, making it the least impactful single heuristic. There is no single heuristic that consistently produces the best gains for all scenarios. For LASER3, XLM-R, and LaBSE, we observe the highest ChrF++ gains of +10.57, +20.61, and +11.40 for CCMatrix-EnTa and CCAligned-EnTa and CCAligned-EnTa corpora, respectively.
2. Impact of Combined Heuristics: The combination of heuristics produced the best score, compared to the best-performing individual heuristic, except for the CCMatrix-SiTa. In this, the combination reduced the dataset size by 54% (Table 15 in Appendix G), resulting in only 98k parallel sentences, which could be the reason for the reduced score.
For CCAligned-SiTa (we exclude CCMatrix due to the dataset size limitation), the gains across multiPLMs are marginal. The same pattern holds with LASER3 across En-Ta and En-Si languages, with the exception of CCMatrix-EnTa corpus. Since LASER3 is already trained with OPUS [^11] https://opus.nlpl.eu/ data, it is believed that LASER3 is already optimized Moon et al. (2023) towards ranking the CCMatrix and CCAligned corpora better. This is further affirmed with the relatively higher NMT scores of the baseline experiments.
dedup+punctNums+(n)gram+sLength+LIDThresh combination appears as the best combination in general. The specific n-value is dependent on the dataset. Only in CCAligned-SiTa with XLM-R, the combination performed best without the LID heuristic. However, we include LIDThresh into the combination for two reasons. (1) Even with LIDThreshold, the score only lags by (-0.19) ChrF++ scores, which is negligible. (2) The most influential individual heuristic is LID-based. In the combination, producing the best gains, the ratio-based heuristic was different based on the dataset/language pair. However, in 80% of cases it had been sentWRatio.
3. Reducing the Disparity Across multiPLMs: In order to calculate the disparity between multiPLMs, we calculated the difference between LASER3 and the other two multiPLMs. This is because LASER3 produced the best result for baseline experiments, and its results consistently improved after applying heuristics.
In Table 7121212Since for the SiTa pair the disparity observed was marginal, we exclude this language-pair from our discussion., the best result of each heuristic category reduces this disparity for all experiments related to CCAlign. Disparity has been reduced on average by 95.87% for CCAligned-XLMR/LaBSE and CCMatrix-LaBSE, compared to LASER3. The disparity between CCMatrix-XLMR EnSi/EnTa models, compared to LASER3 was reduced by 48% only, meaning that XLM-R ranked corpus contains noise which cannot be mitigated by the heuristics alone. However, our hypothesis holds in most cases, and it is safe to say that heuristic-based PDC mitigate the bias brought in by the multiPLM.
Heuristic LASER3 vs XLM-R LASER3 vs LaBSE
Disparity (ChrF++) +/-Disparity wrt Baseline (%) Disparity (ChrF++) +/- Disparity wrt Baseline (%)
CCMatrix
English - Sinhala
Baseline 25.21 16.27
Deduplication - based 18.41 26.97% 8.72 46.40%
Sentence Length - based 24.58 2.50% 2.86 82.42%
LID -based 18.76 25.59% 16.64 -2.27%
Ratio-based 24.10 4.40% 16.00 1.66%
Combined Heuristics 11.92 52.72% 2.16 86.72%
English - Tamil
Disparity 13.16 4.82
Reduction in disparity (dedup) 13.33 -1.29% -0.39 108.09%
Reduction in disparity (sLength) 13.12 0.30% -2.28 147.30%
LID 14.35 -9.04% 4.26 11.62%
Ratio-based 13.74 -4.41% 2.23 53.73%
Combined Heuristics 7.31 44.45% -1.13 123.44%
CCAligned
English - Sinhala
Baseline 12.94 4.76
Deduplication Best 4.91 62.06% 2.50 47.48%
Reduction in disparity (sLength) 5.33 58.81% 1.38 71.01%
LID 2.76 78.67% 2.85 40.13%
Ratio-based 5.42 58.11% 2.80 41.18%
Combined Heuristics 0.60 95.36% 0.59 87.61%
English - Tamil
Disparity 22.73 14.13
Reduction in disparity (dedup) 5.93 73.91% 4.82 65.89%
Reduction in disparity (sLength) 8.87 60.98% 3.46 75.51%
LID 4.62 79.67% 5.23 62.99%
Ratio-based 11.17 50.86% 6.28 55.56%
Combined Heuristics 1.73 92.39% 1.45 89.74%
Table 7: Disparity in ChrF++ points, among the NMT models (XLM-R/LaBSE) with the best scores with respect to LASER3 after applying the individual/combined heuristics. The +/- Disparity is this disparity as a percentage of the baseline disparity.
5.3 Human Evaluation Results
As shown in Table 2, heuristic-based PDC had reduced the noise in the top-ranked samples consistently, irrespective of the language pair and the considered multiPLM. Some gains are quite significant - For example, the amount of correct pairs (C) improvement for CCMatrix-EnSi (LaBSE), CCAligned-EnTa (LaBSE) and CCAligned-EnSi (XLM-R) were all 82%. Before applying the heuristics, the average Error category (E) across language pairs and datasets is 13.50% for LASER3, 36.33% for XLM-R and 72.53% for LaBSE. CS, CCN and UN noise categories were noted to be contributing towards this noise percentage.
However, after heuristic filtration, the error drops drastically. For LASER, XLM-R and LaBSE on average, it is 2.17%, 2.50% and 1.0%, respectively. However, the untranslated (UN) and overlapping text (CCN) were still there. To eliminate these types of noise, we would need to employ an alignment model similar to Steingr´ımsson et al. (2023) or Minh-Cong et al. (2023).
In conclusion, the human evaluation results also indicate that the heuristic-based PDC approach is beneficial for multiPLM ranking in two key ways. First, it produces the top-ranked sentence pairs from multiPLM to be qualitatively comparable. Secondly, it removes the noisy parallel sentences that cause the disparity among the NMT systems trained using ranked corpora based on multiPLMs.
6 NMT Performance based on Training Data Size
We further investigate the effect of the training dataset size on the final NMT performance. We sample 50k, 100k, 150k, and 200k from each ranked corpus, and train NMT models. Results are in Figure 5 of the Appendix H.
For Si-Ta, moderate but consistent improvements were observed from 50K to 100K samples for both CCMatrix and CCAligned ranked corpora, irrespective of the multiPLMs. However, due to the lack of sufficient corpus size, the NMT scores for 150K and 200K were not observable. For En-Ta, NMT scores declined after 100K, while for En-Si, an improvement was observed until 200K, except for the CCMatrix XLM-R ranked sample. Hence, we suspect that noise may be more prominent beyond 100K for En-Ta. In conclusion, increasing corpus size can enhance NMT performance, but only when the underlying data is sufficiently clean and well-aligned.
7 Discussion
We show that applying simple rule-based heuristics to remove noisy sentences before using a multiPLM is better than simply ranking the web-mined corpora based on embeddings derived from multiPLMs as suggested by Ranathunga et al. (2024).
Based on our extensive experiments, dedup+punctNums+(n)gram+sLength+LIDThresh
+sentWRatio is our recommended heuristic combination, based on the justifications provided earlier. Although $n$ can be corpus dependent, we recommend starting with n=5 or 6. In contrast to the recent PDC work Steingr´ımsson (2023); Minh-Cong et al. (2023) that employs several deep learning model-based rigorous filtration pipelines, our technique is much simpler.
8 Conclusion
In this research, we empirically analysed the disparity between the NMT systems trained with the web-mined corpora ranked using embeddings derived from multiPLMs. With a human evaluation, we showed that this disparity is due to different types of noise creeping into the top-ranked portion of corpora when different multiPLMs are used. We made use of rule-based heuristics to remove this noise. After a systematic evaluation of heuristics, we were able to identify optimal heuristic combinations that resulted in higher NMT scores. Therefore, for anyone planning to use web-mined corpora, our recommendation is to first filter out noisy sentences using heuristics and then to do ranking on the embeddings derived from the multiPLM. Human evaluation indicated that even after applying heuristics, some noise remains. In future, we plan to apply techniques such as classification-based approaches to remove such noise, or to apply translation post-editing to improve sentence quality.
9 Limitations and Ethical Concerns
9.1 Limitations
We found that the LID was suboptimal for identifying Si or Ta languages. Therefore, such models would need to be optimised or better LID models would need to be used, such that they are effective in predicting the language label correctly. Due to the lack of funding, we were able to employ only two annotators for the task. However, the results would have been more convincing if the annotation could have been done by three annotators. Furthermore, due to the lack of human annotators, we had to limit this analysis to only three language pairs. We can extend this study to more low-resource language pairs provided that we can find suitable annotators.
9.2 Ethical Concerns
We use publicly available datasets. Due to the dataset size, we did not have the resources or funding to check all the curated parallel sentences manually for offensive content. Fernando et al. (2020) provided the human-crafted dataset to be used in this research. Further, we do not disclose or share any personal details of the annotators publicly or with any other party. We have offered the annotators the standard rate and have settled the payments in full. The only details we disclose are in the Appendix F.
Acknowledgments
Will be provided upon acceptance.
Appendix A Improved Noise Taxonomy
The definitions of all the noise categories in our improved noise taxonomy are given in Table 9. Then, in Table 10, we show examples of parallel sentences which fall under these noise categories.
Table 8: Example parallel sentences which will be separately identified under the new noise category CCN
Error Code: Description
CC: Perfect Translation-pair
Source and target sentences are translation pairs of each other.
CN: Near Perfect Translation-pair
Perfect translation pairs. Just a few spelling, grammar, punctuation, or unnecessary characters have to be handled.
CB: Low-quality Translation-pair
A full sentence or phrase, but a low-quality (boilerplate) translation. Includes under/over translations.
CS: Short Translation Content
Less than 5 words. Translation-wise, correct, but only a short phrase or a few words.
CCN: Overlapping Content
Perfect or near-perfect translation pair, but with overlapping content like numbers, acronyms, or URLs. Sentences longer than 5 words with high overlap.
X: Wrong Translation
Source and target sentences are in the correct languages, but semantically unrelated. Not true translations.
UN: Untranslated Text
The source or target is copied from its counterpart (partial or full). Overlapping untranslated content exceeds 30%. It could have been translated/transliterated.
NL: Not a Language
At least one side is not linguistic content.
WL: Wrong Language
Either the source or the target (or both) is not in the expected language. Up to 30% of acceptable content may be tolerated.
Table 9:  Ranathunga et al. (2024)’s error taxonomy with the CCN category that has been newly added by us
Table 10: Example parallel sentences from the En-Si, En-Ta and Si-Ta, identified during human evaluation.
Appendix B Selection of Languages and Datasets
This section contains details on the selected languages and the web-mined corpora considered under the study.
Sinhala is an Indo-Aryan language spoken primarily in Sri Lanka by the Sinhalese majority. It exhibits complex morphological structures, including rich inflectional and derivational processes, but is classified as a low-resource language due to the scarcity of linguistic resources and tools.
Tamil, a Dravidian language with a rich literary history, is spoken by Tamil communities in Sri Lanka, India, and the global diaspora. Unlike Sinhala, Tamil benefits from a relatively larger digital presence, but it still faces challenges in NLP applications due to morphological complexity, agglutinative grammar, and resource limitations in certain domains.
We obtain the publicly released CCMatrix and CCAligned datasets from the OPUS website [^13] https://opus.nlpl.eu/. Both these datasets support the language pairs, En-Si, En-Ta and Si-Ta, which we consider in this research.
CCMatrix Schwenk et al. (2021) is a web-mined parallel corpus extracted using LASER2-based sentence embeddings to align bitext. While it provides large-scale data, it is highly noisy due to the global mining approach to determine alignments, resulting in misaligned or low-quality translations.
CCAligned El-Kishky et al. (2020) extracts bitext from Common Crawl [^14] https://commoncrawl.org/ using document-level and sentence-level alignment based on multilingual embeddings. Though it improves alignment quality over global bitext-mined corpora, it still contains significant noise, requiring careful filtering for reliable use.
Appendix C Selection of multiPLMs
We include the details on the three multiPLMs considered in this study.
LASER3 Heffernan et al. (2022) (L=12, H=1024, A=4, 250M) [^15] No of Layers, Hidden Layer Dimensions and No of Attention Heads are defined by L, H and A respectively. is a multiPLM favourable for bitext mining and cross-lingual tasks. It improves over previous LASER2 versions by supporting more languages and enhancing alignment quality, but still faces challenges in low-resource settings.
XLM-R Conneau et al. (2020) (L=12, H=768, A=6, 278M parameters) is a transformer-based multiPLM trained on massive amounts of text using masked language modelling. It achieves strong cross-lingual performance but struggles with low-resource languages due to limited training data.
LaBSE Feng et al. (2022) (L=12, H=768, A=12, 471M) is a BERT-based model optimized for multilingual sentence embeddings and bitext retrieval. It provides high-quality cross-lingual representations and is favourable for cross-lingual tasks.
Appendix D Sentence Length Threshold
To determine the optimal threshold for sentence length filtering, we filter sentences less than 3, 4, and 5 words on the source (S) side, target (T) side, and both sides. Similar to other PDC experiments, we then rank the parallel sentence pairs by computing embeddings using three different multiPLMs, and train NMT models using the top 100k ranked data. We conduct these experiments using the CCAligned corpus, for En-Si and En-Ta language-pairs. The ChrF++ results are shown in Table 11. We observe consistently sentence length threshold of 5 yields the best results. Therefore, we select this threshold in all subsequent experiments.
Side LASER3 XLM-R LaBSE
English - Sinhala
Baseline 32.33 19.39 27.57
Sentence Length (flt < 5) S 33.86 26.53 32.97
T 34.88 29.42 33.14
ST 34.83 29.55 33.50
Sentence Length (flt < 4) S 34.25 23.20 31.07
T 33.99 23.06 31.30
ST 34.69 25.64 31.55
Sentence Length (flt < 3) S 34.50 25.94 31.91
T 34.14 26.55 32.49
ST 34.06 27.16 32.76
English - Sinhala
Baseline 40.13 17.40 26.00
Sentence Length (flt < 5) S 41.40 27.60 36.77
T 41.54 30.16 37.61
ST 41.14 32.67 38.08
Sentence Length (flt < 4) S 40.85 24.28 36.03
T 41.15 26.28 37.05
ST 41.52 30.40 37.75
Sentence Length (flt < 3) S 40.90 22.13 33.56
T 41.42 24.37 36.22
ST 40.58 25.45 36.60
Table 11: NMT results in ChrF++ for different sentence length thresholds.
Appendix E NMT Experiments
The experiments were conducted on a NVIDIA Quadro RTX6000 GPU with 24GB VRAM. The hyperparameters used during training, along with the training parameters, are shown in Table 12. We conduct training on the NMT systems for 100 epochs with early stopping criteria and report the results using ChrF++. ChrF++ was chosen over the conventional multi-BLEU Papineni et al. (2002) and sacreBLEU (Post, 2018) because character-level evaluation is more suitable for the considered Sinhala and Tamil languages, which are morphologically rich in nature.
Hyperparameter Argument value
encoder/decoder Layers 6
encoder/decoder attention heads 4
encoder-embed-dim 512
decoder-embed-dim 512
encoder-ffn-embed-dim 2048
decoder-ffn-embed-dim 2048
dropout 0.4
attention-dropout 0.2
optimizer adam
Adam $\beta_{1}$, Adam $\beta_{2}$ 0.9, 0.99
warmup-updates 4000
warmup-init-lr 1e-7
learning rate 1e-3
batch-size 32
patience 6
fp16 True
Table 12: Training parameters for NMT experiments.
Appendix F Human Evaluation
Selection of the Annotators: We select annotators for this task who have translation or machine translation-related experience for a minimum of 2 years and are fluent in both languages of the specific language pair assigned to them. Table 13 shows the years of experience and the qualifications of those annotators who conducted this task.
Annotator Experience Qualification
Annotator 01 22 Diploma in Translation And Interpretation
Annotator 02 9 BSc.(Hons) in Information Technology
Annotator 03 5 MBBS
Annotator 04 4 BA (Hons) in Translation
Annotator 05 3 BSc (Hons) Engineering sp. in Computer Science and Engineering
Annotator 06 2.5 Diploma in Translation And Interpretation
Annotator 07 2.5 BSc Eng (Hons) Electrical & Electronics Engineering
Annotator 08 2.5 BSc (Hons) Engineering sp. in Electrical Engineering
Annotator 09 2 Bachelor of Industrial Information Technology
Table 13: Annotator details with the years of experience and their qualifications.
Resources Provided and Training: All the annotators have had prior experience with a similar task. However, for this annotation work, we provide them with the definitions of the noise categories (Table 9) along with example sentence pairs (Table 10) and the guideline in terms of a flowchart (Figure 3). First, we asked them to do a sample of 30 sentences as a training on the task, and review it. Then, the 1200 sentence pairs to be annotated were shared with each annotator via Google Sheets to be completed.
Compensation: They were paid for each sentence pair they annotated, which was the standard rate for such work.
Appendix G Results Analysis
We show the individual and the combined heuristic which produced the best NMT gains with respect to each multiPLM, dataset and the language pair in the Table 14. The Figure 4 shows the percentage of dedup+ngram experiments exceeding the highest dedup, with respect to the language-pair. Finally, we show the final dataset sizes after applying the heuristic along with the percentage reduction in Table 15.
Heustic/s CCMatrix Heustic/s CCAligned
LASER3 XLM-R LaBSE LASER3 XLM-R LaBSE
Sinhala →Tamil
Baseline 31.08 30.99 31.63 Baseline 35.36 35.97 35.79
Deduplication-based Heuristics
DD+PN+5gram (ST+T) 32.98 32.73 32.60 DD+puntsNums (T) 36.63 36.47 36.86
DD+PN (ST) 35.96 36.71 36.23
DD+PN+7gram 36.73 36.62 36.37
Length-based Heurics
(S) 31.41 31.52 32.30 (T) 36.30 36.71 36.58
(ST) 36.47 35.99 36.60
LID-based Heuristics
LID (S) 31.48 31.36 31.78 LIDThresh (ST) 36.73 36.73 36.80
Ratio-based Heuristics
sentCRatio (ST) 32.28 31.90 32.04 sentWRatio (T) 36.24 36.17 36.46
sentWRatio (ST) 36.44 36.72 36.01
Combined Heuristics
DD+PN+5gram++sentCharRatio 31.45 32.65 31.17 DD+PN+7gram++sentWRatio 36.60 36.85 36.32
DD+PN+5gram++LIDThresh+sentCharRatio 32.64 31.30 32.28 DD+PN+7gram++LIDThresh+sentWRatio0.8 36.83 36.66 37.03
Gain (DD-heuristic vs Baseline) 1.90 1.74 0.97 1.37 0.50 1.07
Gain (sLength vs Baseline) 0.33 0.53 0.67 0.94 0.74 0.79
Gain (LID-heuristic vs Baseline) 0.40 0.37 0.15 1.37 0.76 1.01
Gain (Raio heuristic vs Baseline) 1.20 0.91 0.41 1.08 0.75 0.67
Gain (Overall individual Heuristic vs Baseline) 1.90 1.74 0.97 1.37 0.76 1.07
Gain (Combined Heuristic vs Baseline) 1.56 1.66 0.65 1.47 0.88 1.24
Gain (Combined vs Individual ) -0.34 -0.08 -0.32 0.10 0.12 0.17
English →Sinhala
Baseline 30.76 5.55 14.49 Baseline 32.33 19.39 27.57
Deduplication-based Heuristics
DD+puntsNums (T) 33.89 14.81 26.31 DD+puntsNums+5gram (ST+T) 33.81 30.33 32.74
DD+puntsNums+5gram (T+T) 34.50 16.09 25.78 DD+puntsNums+6gram (ST+T) 35.24 28.21 31.26
Length-based Heurics
sLength (ST) 32.82 8.24 29.96 sLength (T) 34.88 29.42 33.14
sLength (ST) 34.83 29.55 33.50
LID-based Heuristics
LID 31.99 13.32 16.20 LIDThresh (S) 35.73 30.86 32.69
LIDThresh 32.84 14.08 13.71 LIDThresh (ST) 35.11 32.97 32.88
Ratio-based Heuristics
sentWRatio (S) 31.50 7.40 10.86 sentWRatio (S) 34.15 25.97 31.35
sentWRatio (ST) 30.64 7.00 15.50 sentWRatio (ST) 33.85 28.73 31.17
Combined Heuristics
DD+PN+5gram +sLength +LIDThresh +sentWRatio 36.10 23.84 33.94 DD+PN+5gram+sLength+LIDThresh+sentWRatio 36.15 34.50 35.67
DD+PN+5gram+sLength+LIDThresh+sentWRatio>0.8 35.66 24.18 33.19 DD+PN+5gram+sLength+LIDThresh+sentWRatio>0.8 36.26 35.66 35.42
Gain (Dedup-heuristic vs Baseline) 3.74 10.54 11.82 2.91 10.94 5.17
Gain (sLength vs Baseline) 2.06 2.69 15.47 2.55 10.16 5.93
Gain (LID-heuristic vs Baseline) 2.08 8.53 1.71 3.40 13.58 5.31
Gain (Raio heuristic vs Baseline) 0.74 1.85 1.01 1.82 9.34 3.78
Gain (Overall individual Heuristic vs Baseline) 3.74 10.54 15.47 3.40 13.58 5.93
Gain (Combined Heuristic vs Baseline) 5.34 18.63 19.45 3.93 16.27 8.10
Gain (Combined vs Individual ) 1.60 8.09 3.98 0.53 2.69 2.17
English →Tamil
Baseline 19.02 5.86 14.20 Baseline 40.13 17.40 26.00
Deduplication-based Heuristics
DD-7gram (T) 18.15 7.37 21.96 DD+puntsNums+4gram (ST+T) 41.82 35.90 37.08
DD+puntNums (T) 21.57 8.24 20.41 DD+puntsNums+6gram (ST+ST) 41.90 35.97 35.94
Length-based Heurics
sLength (T) 18.52 6.33 21.73 sLength (T) 41.54 30.16 37.61
sLength (ST) 19.45 5.33 20.79 sLength (ST) 41.14 32.67 38.08
LID-based Heuristics
LIDThresh (T) 29.59 15.24 24.51 LIDThresh ST) 42.63 38.01 37.40
LIDThresh (ST) 28.93 15.16 25.33
Ratio-based Heuristics
STRatio 20.52 5.40 18.29 sentWRatio (S) 42.05 29.70 35.53
sentCRatio (T) 19.90 6.78 12.51 sentWRatio (ST) 41.05 30.88 35.77
Combined Heuristics
DD+PN+7gram+slength+LIDThresh+STRatio 30.67 23.36 31.80 DD+PN+6gram+sLength+LIDThresh+sentWRatio 43.47 41.74 41.06
DD+PN+6gram+sLength+LIDThresh+sentWRatio > 0.8 42.08 40.56 42.02
Gain (Dedup-heuristic vs Baseline) 2.55 2.38 7.76 1.77 18.57 11.08
Gain (sLength vs Baseline) 0.43 0.47 7.53 1.41 15.27 12.08
Gain (LID-heuristic vs Baseline) 10.57 0.92 10.31 2.50 20.61 11.40
Gain (Raio heuristic vs Baseline) 1.50 0.92 4.09 1.92 13.48 9.77
Gain (Overall individual Heuristic vs Baseline) 10.57 2.38 10.31 2.50 20.61 12.08
Gain (Combined Heuristic vs Baseline) 11.65 17.50 17.60 1.92 12.30 16.02
Gain (Combined vs Individual ) 1.08 8.12 6.47 0.84 3.73 3.94
Table 14: Shows the best NMT performance produced using the individual heuristics as well as combinations of heuristics. The values in bold indicate the highest NMT score obtained for a given heuristic or heuristic combination. The values underlined are the highest among the individual heuristics. Highlighted in green are the overall best values. Here DD+PN is Deduplication+punctNums, SL is sLength and LT is LIDThresh.
Heuristic Applicable Side Sinhala - Tamil English - Sinhala English - Tamil
CCMatrix CCAligned CCMatrix CCAligned CCMatrix CCAligned
Dataset Size % Reduction Dataset Size % Reduction Dataset Size % Reduction Dataset Size % Reduction Dataset Size % Reduction Dataset Size % Reduction
Baseline 215965 260119 6270800 619730 7291118 880568
$D$ S 189654 250038 3.88% 6146819 1.98% 570768 7.90% 6378607 12.52% 797071 9.48%
T 209461 247176 4.98% 3242950 48.28% 562088 9.30% 4754106 34.80% 780355 11.38%
ST 183904 243384 6.43% 3176145 49.35% 537581 13.26% 4060447 44.31% 736212 16.39%
DD-4gram S 176590 18.23% 189218 27.26% 4171884 33.47% 403993 34.81% 4802654 34.13% 440248 50.00%
T 196538 9.00% 207603 20.19% 2751819 56.12% 423752 31.62% 4271550 41.41% 549621 37.58%
ST 172440 20.15% 184159 29.20% 2035282 67.54% 355790 42.59% 100.00% 365648 58.48%
DD-5gram S 188457 12.74% 217733 16.29% 4486108 28.46% 481021 22.38% 5104643 29.99% 558504 36.57%
T 204045 5.52% 226738 12.83% 3071693 51.02% 499307 19.43% 4516819 38.05% 638600 27.48%
ST 185915 13.91% 216389 16.81% 2374578 62.13% 446838 27.90% 3319964 54.47% 487465 44.64%
DD-6gram S 196194 9.15% 232604 10.58% 5383674 14.15% 528525 14.72% 5309083 27.18% 639961 27.32%
T 200310 7.25% 237467 8.71% 3142124 49.89% 539513 12.94% 4590987 37.03% 681186 22.64%
ST 196194 9.15% 233792 10.12% 2859756 54.40% 505429 18.44% 3489629 52.14% 570403 35.22%
DD-7gram S 200899 6.98% 240704 7.46% 5701801 9.07% 554025 10.60% 5718913 21.56% 679750 22.81%
T 204485 5.32% 244007 6.19% 3170538 49.44% 561285 9.43% 4631486 36.48% 703950 20.06%
ST 200898 6.98% 246104 5.39% 3021457 51.82% 538898 13.04% 3745771 48.63% 611821 30.52%
DD+N S 182386 15.55% 260119 0.00% 6105433 2.64% 505828 18.38% 6337285 13.08% 683882 22.34%
T 201551 6.67% 218980 15.82% 3225979 48.56% 502971 18.84% 4722644 35.23% 675627 23.27%
ST 176040 18.49% 216238 16.87% 3158067 49.64% 476379 23.13% 4031459 44.71% 631485 28.29%
DD+PN S 180380 16.48% 215100 17.31% 5931349 5.41% 494778 20.16% 6194331 15.04% 668849 24.04%
T 198352 8.16% 212341 18.37% 3197186 49.01% 492801 20.48% 4666158 36.00% 660902 24.95%
ST 173804 19.52% 207197 20.35% 3130297 50.08% 465617 24.87% 3987832 45.31% 616702 29.97%
DD+PN+4gram ST+T 289248 53.33%
DD+PN+5gram ST + T 167022 22.66% 187250 28.01% 3044520 51.45% 380146 38.66%
DD+PN+6gram ST + T 169784 21.38% 189060 27.32% 428939 30.79% 4547759 37.63% 464424 47.26%
DD+PN+7gram T +T 196221 24.56% 4620008 36.64%
SL S 150094.00 30.50% 188061 27.70% 5088747 18.85% 411474.00 33.60% 6498956 10.86% 595057 32.42%
T 100799.00 53.33% 161363 37.97% 3670963 41.46% 377708.00 39.05% 4267495 41.47% 517516 41.23%
ST 96264.00 55.43% 157978 39.27% 3341564 46.71% 348829.00 43.71% 4134919 43.29% 491207 44.22%
LID S 192377.00 10.92% 241617 7.11% 6200355 1.12% 479589.00 22.61% 7210848 1.10% 669260 24.00%
T 186720.00 13.54% 241863 7.02% 6066681 3.26% 575298.00 7.17% 6800923 6.72% 794143 9.81%
ST 178276.00 17.45% 231619 10.96% 6010065 4.16% 457639.00 26.16% 6743988 7.50% 625281 28.99%
LT S 181470.00 15.97% 227791 12.43% 6120792 2.39% 398272.00 35.73% 6120793 16.05% 564870 35.85%
T 172726.00 20.02% 222290 14.54% 5990169 4.48% 546472.00 11.82% 5990170 17.84% 731484 16.93%
ST 162777.00 24.63% 208644 19.79% 5877142 6.28% 377579.00 39.07% 5877143 19.39% 518010 41.17%
STRatio - 170168.00 21.21% 229101 11.92% 4293239 31.54% 459473.00 25.86% 4051888 44.43% 679820 22.80%
sentWRatio S 199788 7.49% 246908 5.08% 6232528 0.61% 546460 11.82% 7231531 0.82% 743624 15.55%
T 196812 8.87% 250151 3.83% 6198124 1.16% 552798 10.80% 7176111 1.58% 745221 15.37%
ST 193989 10.18% 245161 5.75% 6177212 1.49% 531963 14.16% 7138854 2.09% 717159 18.56%
sentCRatio S 212287 1.70% 224031 13.87% 6262297 0.14% 594169 4.12% 7252444 0.53% 832611 5.45%
T 212877 1.43% 218726 15.91% 6261991 0.14% 596346 3.77% 7281050 0.14% 851343 3.32%
ST 211661 1.99% 215151 17.29% 6257079 0.22% 588310 5.07% 7247298 0.60% 826866 6.10%
Combined Heuristics
DD+PN+ngram (SiTa-CCMatrix n=5, SiTa-CCAligned n= 7 EnSi-CCMatrix/CCAligned n=5, EnTa-CCMatrix n=7, EnTa-CCAligned n=6)
+SL T+ST 117198 45.73% 143919 44.67% 2245307 64.19% 240086 61.26% 3462458 52.51% 321003 63.55%
+LT T+ST 130831 39.42% 162543 37.51% 2880530 54.06% 239144 61.41% 2876818 60.54% 306352 65.21%
+sentWRatio T+S 154028 28.68% 170715 34.37% 2993730 52.26% 337634 45.52% 3377988 53.67% 427587 51.44%
+SL+LT T+ ST 99207 54.06% 127284 51.07% 2200195 64.91% 180731 70.84% 4330140 40.61% 241679 72.55%
+SL+sentWRatio T+ST+ST 116344 46.13% 127035 51.16% 2241513 64.25% 224542 63.77% 2794637 61.67% 298141 66.14%
+SL+LT+sentWRatio T+ST+ST+S 127188 41.11% 117970 54.65% 2197629 64.95% 177711 71.32% 2726087 62.61% 237105 73.07%
+SL+LT+sentWRatio>0.8 T+ST+ST+ST 179984 16.66% 99311 61.82% 2149037 65.73% 161869 73.88% 2203591 69.78% 214639 75.62%
+SL+LT+sentCRatio T+ST+ST+ST 98894 54.21% NA NA NA NA NA
+SL+LT+STRatio T+ST+ST+STR 82866 61.63% NA NA NA 2129744 70.79% NA
Table 15: Shows the final corpus sizes after applying heuristics along with the reduction percentage. Here DD+PN is Deduplication+punctNums, SL is sLength and LT is LIDThresh. NA corresponds to the experiments that are not applicable for the language-pair.
Figure 3: Shows the annotation guideline document in terms of a flow chart. This shows the priority of the noise category to be selected prior to declaring the annotation class.
Figure 4: Percentage of dedup+ngram experiments exceeding the best result of dedup with respect to the Language-pair.
Appendix H NMT Performance on Training Dataset
Figure 5 shows the variation in NMT results by varying the dataset size with 50k, 100, 150k and 200k for each language-pair.
(a) En-Si
(b) En-Ta
(c) Si-Ta
Figure 5: ChrF++ scores of NMT systems trained by varying the training dataset size.

### v3

@@
 Ranking sentence pairs using similarity scores on sentence embeddings derived from Pre-trained Multilingual Language Models (multiPLMs) is the most common PDC technique.
-However, previous research has shown that the choice of the multiPLM significantly impacts the quality of the filtered parallel corpus, and the NMT models trained using such data show a disparity across multiPLMs.
+However, previous research has shown that the choice of the multiPLM significantly impacts the quality of the filtered parallel corpus, and the Neural Machine Translation (NMT) models trained using such data show a disparity across multiPLMs.
 This paper shows that this disparity is due to different multiPLMs being biased towards certain types of sentence pairs, which are treated as noise from an NMT point of view.
 We show that such noisy parallel sentences can be removed to a certain extent by employing a series of heuristics.
-The curated corpus leads to producing better results by NMT models while minimizing the disparity across multiPLMs.
-We publicly release the source code and the curated datasets [^1] Source codes and Datasets will be released upon acceptance.
+The NMT models, trained using the curated corpus, lead to producing better results while minimizing the disparity across multiPLMs.
+We publicly release the source code and the curated datasets [^1] https://github.com/aloka-fernando/Heuristic_based_parallel_corpus_filtration.
 Improving the Quality of Web-mined Parallel Corpora of Low-Resource Languages using Debiasing Heuristics
 Aloka Fernando1, Nisansa de Silva1, Menan Velayuthan1, Charitha Rathnayake1 Surangika Ranathunga 2 {alokaf,NisansaDdS,velayuthan.22,charitha.18}@cse.mrt.ac.lk s.ranathunga@massey.ac.nz 1Dept.
-of Computer Science and Engineering, University of Moratuwa,10400, Sri Lanka 2Massey University, Palmerston North, New Zealand, 4443 Correspondence: alokaf@cse.mrt.ac.lk
+of Computer Science and Engineering, University of Moratuwa,10400, Sri Lanka 2School of Mathematical and Computational Sciences, Massey University, New Zealand Correspondence: alokaf@cse.mrt.ac.lk
 1 Introduction
@@
 (2020) are examples of such web-mined corpora, which cover Low-Resource Languages (LRLs) as well.
-However, quality audits of these corpora, as shown in Table 1, reveal substantial noise, which negatively impacts the NMT models Khayrallah and Koehn (2018).
-Noise Category A B C D E
+As summarised in Table 1, quality audits of these corpora reveal different types of noise.
+Consequently, training NMT models on such noisy parallel data leads to low-quality translations Khayrallah and Koehn (2018).
+Parallel Sentence Quality Category A B C D E
 Perfect translations - - - Y Y
@@
 Short Sentences Y - Y - Y
-Mismatch numbers Y - - -
+Mismatch numbers - Y - - -
 Machine Translated Sentences - - Y Y -
@@
 Not a Language Y Y Y Y Y
-Table 1: Noise types in parallel corpora, as identified by Khayrallah and Koehn (2018) (A), Bane et al.
+Table 1: Parallel sentence quality categories used in quality audits by Khayrallah and Koehn (2018) (A), Bane et al.
 (2022) (B), Herold et al.
@@
 (2024) (E).
-Parallel Data Curation (PDC) aims at improving the quality of such web-mined parallel corpora by filtering high-quality parallel sentence pairs.
-The importance of PDC for LRLs has been emphasised with the introduction of WMT shared tasks Sloto et al.
+Parallel Data Curation (PDC) aims at extracting high-quality parallel sentence pairs from noisy web-mined corpora.
+The importance of PDC for LRLs has been emphasised with the introduction of PDC shared tasks Sloto et al.
 (2023); Koehn et al.
@@
 (2019), recent PDC techniques follow a scoring and ranking mechanism using embeddings obtained from a Multilingual Language Model (multiPLM).
-Then, a subset of the top-ranked sentence pairs is selected to train the NMT model.
-However, the quality of these top-ranked pairs depends on the selected multiPLM (Ranathunga et al., 2024; Moon et al., 2023).
-To investigate this further, we obtain embeddings from three multiPLMs: LASER3 Heffernan et al.
+During the scoring step, the semantic similarity is calculated between the source and target sentence embeddings for each sentence pair.
+Then the sentence pairs are ranked in descending order of the similarity score.
+Finally, a subset of the top-ranked sentence pairs is selected to train the NMT model.
+However, the quality of these top-ranked pairs depends on the chosen multiPLM (Ranathunga et al., 2024; Moon et al., 2023).
+To investigate the impact of using different multiPLMs on the PDC task, we conduct an initial analysis.
+We obtain embeddings from three multiPLMs: LASER3 Heffernan et al.
 (2022), XLM-R Conneau et al.
@@
 (2020) datasets and rank them in descending order.
-Then we train NMT models (Section 4.4) using the top-ranked 100k sentence pairs from each corpus.
-Experiments are carried out for English-Sinhala (En-Si), English-Tamil (En-Ta) and Sinhala(Si-Ta).
-As shown in Figure 1, there is a significant disparity among the results, mainly for En-Si and En-Ta language pairs.
-Figure 1: Baseline NMT scores using top-ranked sentence pairs from CCMatrix and CCAligned, using embeddings obtained from LASER3, XLM-R, and LaBSE.
-A manual inspection of the En-Si top-ranked 1000 sentences reveal that different multiPLMs prioritise different sentence characteristics when ranking parallel sentences.
-For example, sentence pairs ranked top by XLM-R are mostly short and contain overlapping text such as numbers, acronyms, and URLs.
-LaBSE also favours sentences with numbers and date overlaps, while LASER3 prioratizes relatively better full length sentence-pairs.
-The impact of the nature of the top-ranked sentences is evident in Figure 1, where results related to LASER3 outperform its two counterparts by a significant margin, in most of the experiments.
-In order to carry out a more quantitative evaluation, we randomly selected 100 sentence pairs from the top 1000 sentences from the aforementioned ranked corpora and carried out a human evaluation.
+Then we train NMT models (Section 4.4) using the top-ranked 100k sentence pairs from each corpus and report the ChrF++ scores.
+Experiments are carried out for English-Sinhala (En-Si), English-Tamil (En-Ta) and Sinhala-Tamil (Si-Ta) language-pairs.
+Sinhala, Tamil, and English belong to three distinct linguistic groups: Indo-Aryan, Dravidian, and Germanic (respectively).
+As shown in Figure 1, there is a significant difference, i.e.
+a disparity among these results, mainly for En-Si and En-Ta language pairs.
+Figure 1: Baseline NMT scores using top-ranked sentence pairs from CCMatrix and CCAligned corpora, with embeddings from LASER3, XLM-R, and LaBSE.
+A random inspection of the En-Si top-ranked sentences reveals that different multiPLMs prioritise different sentence characteristics when ranking parallel sentences.
+For example, sentence pairs ranked top with XLM-R embeddings are mostly short and have high overlapping text such as numbers, acronyms, and URLs.
+LaBSE embeddings also result in sentences with numbers and date overlaps, while LASER3 embeddings result in relatively better full-length sentence pairs.
+This observation sheds light on the disparity in NMT results shown in Figure 1 - the sentences ranked top when using XLM-R and LaBSE may have less linguistic content for NMT models to learn the translation.
+In order to carry out a more systematic evaluation, we randomly select 100 sentence pairs from the top 100K parallel sentences from the aforementioned ranked corpora and carried out a human evaluation.
 Ranathunga et al.
-(2024)’s is the most comprehensive error taxonomy available for this task.
-However, to better capture the noise identified during the manual observation, we extended this error taxonomy (Section 3).
-Human evaluation results in Table 2 show that compared to LASER3, the error category percentage (E) is excessively high for XLM-R and LaBSE.
-The most contributing types of errors are untranslated text (UN), short sentences (CS) and sentences with number/acronym/URL overlaps (CCN).
-Examples for these are shown in Table 10 in Appendix F.
+(2024)’s is the most comprehensive taxonomy available for this task.
+To better capture the noise observed during human evaluation, we extend this taxonomy (Section 3).
+Human evaluation results (Table 2) indicate that the mean noise percentages are 98%, 95%, and 70% when the corpora are ranked using LaBSE, XLM-R, and LASER3 embeddings, respectively.
+Further, the noise types in the categories of untranslated text (UN), short sentences (CS) and sentences with high-overlapping non-translatable text (CCN) could be observed, contributing to the reported percentages.
+Examples for these noise types are shown in Table 10 in Appendix A.
+We provide a detailed discussion of these results in Section 5.3.
+These findings suggest that inherent biases in multiPLMs lead to noisy parallel sentences being ranked highly, which in turn contributes to disparities across NMT models.
 $C$ CN CB     C     CS CCN UN X WL NL     E     $C$ CN CB     C     CS CCN UN X WL NL     E     $C$ CN CB     C     CS CCN UN X WL NL     E
@@
 CCMatrix
-LASER3-Before 8.00% 27.00% 2.00%     37.00%     14.00% 14.00% 34.00% 1.00% 0.00% 0.00%     63.00%     17.00% 7.00% 4.00%     28.00%     7.00% 10.00% 55.00% 0.00% 0.00% 0.00%     72.00%     0.00% 3.00% 2.00%     5.00%     0.00% 0.00% 95.00% 0.00% 0.00% 0.00%     95.00%
-LASER3-After 16.00% 56.00% 1.00%     73.00%     13.00% 4.00% 10.00% 0.00% 0.00% 0.00%     27.00%     39.00% 39.00% 7.00%     85.00%     0.00% 7.00% 8.00% 0.00% 0.00% 0.00%     15.00%     6.00% 61.00% 20.00%     87.00%     0.00% 3.00% 10.00% 0.00% 0.00% 0.00%     13.00%
-XLM-R-Before 1.00% 10.00% 0.00%     11.00%     40.00% 19.00% 29.00% 0.00% 1.00% 0.00%     89.00%     1.00% 0.00% 0.00%     1.00%     13.00% 4.00% 80.00% 2.00% 0.00% 0.00%     99.00%     0.00% 0.00% 2.00%     2.00%     3.00% 5.00% 90.00% 0.00% 0.00% 0.00%     98.00%
-XLM-R - After 0.00% 20.00% 2.00%     22.00%     13.00% 29.00% 35.00% 0.00% 1.00% 0.00%     78.00%     3.00% 8.00% 26.00%     37.00%     0.00% 2.00% 53.00% 8.00% 0.00% 0.00%     63.00%     0.00% 39.00% 31.00%     70.00%     1.00% 3.00% 21.00% 4.00% 0.00% 1.00%     30.00%
-LaBSE - Before 4.00% 6.00% 0.00%     10.00%     74.00% 7.00% 9.00% 0.00% 0.00% 0.00%     90.00%     13.00% 2.00% 0.00%     15.00%     63.00% 14.00% 8.00% 0.00% 0.00% 0.00%     85.00%     0.00% 9.00% 2.00%     11.00%     34.00% 7.00% 48.00% 0.00% 0.00% 0.00%     89.00%
-LaBSE - After 29.00% 33.00% 0.00%     62.00%     2.00% 32.00% 4.00% 0.00% 0.00% 0.00%     38.00%     87.00% 7.00% 3.00%     97.00%     0.00% 1.00% 2.00% 0.00% 0.00% 0.00%     3.00%     36.00% 53.00% 4.00%     93.00%     1.00% 3.00% 2.00% 1.00% 0.00% 0.00%     7.00%
+LASER3 BF 8% 27% 2%     37%     14% 14% 34% 1% 0% 0%     63%     17% 7% 4%     28%     7% 10% 55% 0% 0% 0%     72%     0% 3% 2%     5%     0% 0% 95% 0% 0% 0%     95%
+AF 16% 56% 1%     73%     13% 4% 10% 0% 0% 0%     27%     39% 39% 7%     85%     0% 7% 8% 0% 0% 0%     15%     6% 61% 20%     87%     0% 3% 10% 0% 0% 0%     13%
+XLM-R BF 1% 10% 0%     11%     40% 19% 29% 0% 1% 0%     89%     1% 0% 0%     1%     13% 4% 80% 2% 0% 0%     99%     0% 0% 2%     2%     3% 5% 90% 0% 0% 0%     98%
+AF 0% 20% 2%     22%     13% 29% 35% 0% 1% 0%     78%     3% 8% 26%     37%     0% 2% 53% 8% 0% 0%     63%     0% 39% 31%     70%     1% 3% 21% 4% 0% 1%     30%
+LaBSE BF 4% 6% 0%     10%     74% 7% 9% 0% 0% 0%     90%     13% 2% 0%     15%     63% 14% 8% 0% 0% 0%     85%     0% 9% 2%     11%     34% 7% 48% 0% 0% 0%     89%
+AF 29% 33% 0%     62%     2% 32% 4% 0% 0% 0%     38%     87% 7% 3%     97%     0% 1% 2% 0% 0% 0%     3%     36% 53% 4%     93%     1% 3% 2% 1% 0% 0%     7%
 CCAligned
-LASER3-Before 3.00% 24.00% 3.00%     30.00%     34.00% 19.00% 17.00% 0.00% 0.00% 0.00%     70.00%     2.00% 22.00% 8.00%     32.00%     13.00% 30.00% 23.00% 2.00% 0.00% 0.00%     68.00%     2.00% 23.00% 18.00%     43.00%     13.00% 27.00% 17.00% 0.00% 0.00% 0.00%     57.00%
-LASER3-After 5.00% 79.00% 2.00%     86.00%     0.00% 9.00% 4.00% 1.00% 0.00% 0.00%     14.00%     13.00% 58.00% 14.00%     85.00%     0.00% 0.00% 13.00% 2.00% 0.00% 0.00%     15.00%     3.00% 67.00% 10.00%     80.00%     0.00% 8.00% 12.00% 0.00% 0.00% 0.00%     20.00%
-XLM-R-Before 0.00% 0.00% 2.00%     2.00%     48.00% 49.00% 0.00% 0.00% 0.00% 1.00%     98.00%     2.00% 0.00% 0.00%     2.00%     72.00% 20.00% 6.00% 0.00% 0.00% 0.00%     98.00%     0.00% 8.00% 4.00%     12.00%     42.00% 16.00% 15.00% 8.00% 0.00% 7.00%     88.00%
-XLM-R - After 20.00% 33.00% 4.00%     57.00%     1.00% 22.00% 19.00% 0.00% 1.00% 0.00%     43.00%     18.00% 18.00% 20.00%     56.00%     0.00% 6.00% 34.00% 4.00% 0.00% 0.00%     44.00%     6.00% 46.00% 30.00%     82.00%     0.00% 9.00% 9.00% 0.00% 0.00% 0.00%     18.00%
-LaBSE - Before 0.00% 1.00% 0.00%     1.00%     69.00% 26.00% 3.00% 0.00% 0.00% 1.00%     99.00%     0.00% 1.00% 0.00%     1.00%     97.00% 2.00% 0.00% 0.00% 0.00% 0.00%     99.00%     0.00% 1.00% 0.00%     1.00%     97.00% 0.00% 0.00% 0.00% 0.00% 2.00%     99.00%
-LaBSE - After 15.00% 34.00% 0.00%     49.00%     2.00% 43.00% 6.00% 0.00% 0.00% 0.00%     51.00%     45.00% 27.00% 3.00%     75.00%     1.00% 19.00% 5.00% 0.00% 0.00% 0.00%     25.00%     19.00% 45.00% 3.00%     67.00%     0.00% 22.00% 11.00% 0.00% 0.00% 0.00%     33.00%
-Table 2: Human evaluation results for CCMatrix and CCAligned for En-Si, En-Ta and Si-Ta.
-Results are reported for LASER3, XLM-R, and LaBSE before and after applying heuristics.
+LASER3 BF 3% 24% 3%     30%     34% 19% 17% 0% 0% 0%     70%     2% 22% 8%     32%     13% 30% 23% 2% 0% 0%     68%     2% 23% 18%     43%     13% 27% 17% 0% 0% 0%     57%
+AF 5% 79% 2%     86%     0% 9% 4% 1% 0% 0%     14%     13% 58% 14%     85%     0% 0% 13% 2% 0% 0%     15%     3% 67% 10%     80%     0% 8% 12% 0% 0% 0%     20%
+XLM-R BF 0% 0% 2%     2%     48% 49% 0% 0% 0% 1%     98%     2% 0% 0%     2%     72% 20% 6% 0% 0% 0%     98%     0% 8% 4%     12%     42% 16% 15% 8% 0% 7%     88%
+AF 20% 33% 4%     57%     1% 22% 19% 0% 1% 0%     43%     18% 18% 20%     56%     0% 6% 34% 4% 0% 0%     44%     6% 46% 30%     82%     0% 9% 9% 0% 0% 0%     18%
+LaBSE BF 0% 1% 0%     1%     69% 26% 3% 0% 0% 1%     99%     0% 1% 0%     1%     97% 2% 0% 0% 0% 0%     99%     0% 1% 0%     1%     97% 0% 0% 0% 0% 2%     99%
+AF 15% 34% 0%     49%     2% 43% 6% 0% 0% 0%     51%     45% 27% 3%     75%     1% 19% 5% 0% 0% 0%     25%     19% 45% 3%     67%     0% 22% 11% 0% 0% 0%     33%
+Table 2: The Human evaluation results showing the average percentage for each annotation class for CCMatrix and CCAligned corpora for the En-Si, En-Ta and Si-Ta language pairs.
+The sample sentences have been obtained before (BF) and after (AF) applying the heuristics.
 (C) - overall correct percentage considering $C$ (perfect translation), CN (near perfect) and CB (boilerplate).
-(E) - overall error percentage considering CCN (correct with overlaps), CS (correct but short sentence), X (wrong translation), UN (untranslated), WL (wrong language), NL (not a language).
-We hypothesis that at least some of these noisy sentences can be filtered using rule-based heuristics.
+(E) - overall error percentage considering CCN (Non-translatable overlaps), CS (correct but short sentence), X (wrong translation), UN (untranslated), WL (wrong language), NL (not a language).
+We hypothesise that some of these noisy sentences can be filtered using rule-based heuristics.
 Although applying heuristics is a common approach to improving the quality of parallel corpora Sloto et al.
-(2023); Steingr´ımsson et al.
-(2023), the use of heuristics has not been consistent.
-They have limited their analysis to a single commonly used heuristic, or have varied the criteria or thresholds of the heuristics arbitrarily from one study to another.
-Further, their work does not explore the potential of combining multiple heuristics.
+(2023); Steingrímsson et al.
+(2023), the use of heuristics has not been consistent in the PDC tasks Steingrimsson (2023); Bala Das et al.
+(2023); Aulamo et al.
+(2020).
+Previous research either applied a single heuristic or a subset of commonly used heuristics during pre-processing, with threshold choices varying across studies.
+Further, they did not systematically analyse the impact of heuristic-based filtering on the NMT performance.
 In this research, we incorporate heuristics proposed in previous studies, along with a new heuristic of our own, and conduct an empirical study to investigate whether a more refined selection of top-ranked parallel sentences can be identified by systematically combining these heuristics.
@@
   • We extend  Ranathunga et al.
-(2024)’s error taxonomy with a new error category to capture additional type of noise, which refers to sentence pairs with high-overlapping instances of numbers, acronyms, URLs.
-  • We conduct a systematic study to analyse the impact of rule-based heuristics in filtering noisy sentences from the web-mined corpora and identify an optimal combination of heuristics that works across the corpora and languages considered in the study.
+(2024)’s parallel sentence categorization taxonomy with a new error category to capture an additional type of noise, which refers to sentence pairs with high-overlapping non-translatable text such as numbers, acronyms, URLs, etc.
   • We empirically show that applying heuristics before ranking sentences based on embeddings derived from multiPLMs results in higher NMT scores, and reduces the disparity across multiPLMs.
+  • We conduct a systematic study to analyse the impact of rule-based heuristics in filtering noisy sentences from the web-mined corpora and identify an optimal combination of heuristics that works across corpora and languages considered in the study.
   • We conduct a human evaluation to assess the impact of noise filtering across three multiPLMs.
@@
 2.1 MultiPLMs for PDC
-While employing a multiPLM for PDC is common, past research experimented with only one multiPLM at a time.
-For example, in the WMT2023 shared task, LASER2 was utilized to set the task baseline, whereas Steingr´ımsson (2023) and Gala et al.
-(2023) only used LaBSE.
-Therefore, the disparities across multiPLMs and biases specific to each multiPLM have not come into light.
+While employing a multiPLM for PDC is common, existing research experimented with only one multiPLM at a time.
+For example, in the WMT2023 shared task Sloto et al.
+(2023), LASER2 was utilised to set the task baseline, whereas for the same task, Steingrimsson (2023) used LaBSE.
+Gala et al.
+(2023) also used LaBSE for their work.
+Therefore, the disparities across multiPLMs and biases specific to each multiPLM have not come to light.
 On the other hand, studies conducted by Ranathunga et al.
@@
 (2023) observe that this is due to biases in multiPLMs, which tend to rank noisy parallel sentences highly.
-However, there has been no study to identify these biases.
+However, there has been no systematic study to identify these biases.
 2.2 Identifying Noise in Web-mined Corpora
@@
 Ranathunga et al.
-(2024)’s taxonomy (See Table 9 in Appendix F) is an extension of Kreutzer et al.
+(2024)’s taxonomy (See Table 9 in Appendix A) is an extension of Kreutzer et al.
 (2022)’s taxonomy.
-They first ranked the datasets based on embeddings from a multiPLM, and then selected random samples from top and bottom portions.
-Their human evaluation based on the taxonomy reported that the quality of the parallel sentences varies heavily depending on the selected portion.
+They first ranked the datasets based on embeddings from a multiPLM, and then selected random samples from the top and bottom portions and conducted a quality audit.
+Their human evaluation reported that the quality of the parallel sentences varies heavily depending on the selected portion.
+These studies primarily focus on identifying general noise types; however, their effectiveness in quantifying the specific noise types to which multiPLMs are biased has not yet been evaluated to the best of our knowledge.
 2.3 Heuristic-based PDC
-Commonly used rule-based heuristics can be categorised into four groups as described below:
-De-duplication based (Dedup): Removing identical duplicates in the monolingual sides is a common practice Costa-jussà et al.
-(2022), while De-duplicating after removing non-alpha characters and punctuations Bala Das et al.
+In existing work, the commonly used rule-based heuristics can be categorised into four groups as described below:
+Deduplication-based (Dedup): Removing identical duplicates from the monolingual sides is a common practice Costa-jussà et al.
+(2022).
+Additionally deduplicating after removing non-alpha characters and punctuations Bala Das et al.
 (2023) could be found as its variants.
+While this step has been applied during the pre-processing stage, an empirical study has not been conducted to evaluate its impact on the final NMT performance.
 Length-based (sLength) : Gala et al.
 (2023) and Aulamo et al.
-(2023) have identified the removal of short sentences as a potential heuristic.
-Short sentences hinder NMT models in two ways Koehn and Knowles (2017): by providing insufficient syntactic and semantic information, or resulting in an overfitting situation.
-LID-based (LID): Language Identification is used to remove fully/partially untranslated text and content in the wrong language Steingr´ımsson et al. (2023); Gala et al. (2023); Zhang et al. (2020).
+(2023) have removed short sentences as a potential heuristic.
+Short sentences hinder NMT models in two ways Koehn and Knowles (2017): by providing insufficient syntactic and semantic information, and can result in an overfitting situation.
+LID-based (LID): Language Identification is used to remove fully/partially untranslated text and content in the wrong language Steingrímsson et al. (2023); Gala et al. (2023); Zhang et al. (2020).
 Ratio-based : Ratio-based heuristics identify and remove sentences that show significant structural imbalances between the source and target sentences.
 It is based on the assumption that, well-aligned sentence pairs tend to maintain consistent ratios in terms of character count, word count, or token distribution.
-We observe three common types of ratio-based heuristics: (1) source-to-target sentence length ratio (STRatio(Rossenbach et al., 2018; Gale and Church, 1993), (2) alpha-only words to sentence words ratio (sentWRatio(Aulamo et al., 2020) and (3) alpha-only character ratio with respect to the sentence characters (sentCRatioHangya and Fraser (2018).
-However, the impact of these heuristics in isolation and as a combination had not been evaluated systematically.
+We observe three common types of ratio-based heuristics: (1) source-to-target sentence length ratio (STRatio(Rossenbach et al., 2018; Gale and Church, 1993), (2) alpha-only words to sentence words ratio (sentWRatio(Velayuthan et al., 2024; Aulamo et al., 2020) and (3) alpha-only character ratio with respect to the sentence characters (sentCRatioHangya and Fraser (2018).
+However, the impact of these heuristics in isolation and as a combination had not been evaluated systematically in the context of NMT.
 Ranathunga et al. (2024)
@@
 Revision
-CCN Perfect/near perfect translation where more than 30% of the overlapping content is numbers/acronyms/URLs/etc
+- CCN Perfect/near perfect translation where more than 30% of the overlapping content is non-translatable such as numbers/acronyms/URLs/email etc.
 Short Sentences (Max 3 words) CS Less than 5 words on either side
 Wrong Language WL Specifically set a threshold as 30%
-Table 3: A comparison of the improved taxonomy against Ranathunga et al. (2024)’s (only showing the changes).
+Table 3: A comparison of the improved taxonomy against Ranathunga et al. (2024)’s (only showing the changes). See Table 9 in Appendix A for the full taxonomy.
 3 Methodology
-Improved Taxonomy for Noise Although translation pairs can have overlapping URLs, acronyms, etc, excessive inclusions of such content in a sentence (e.g.
-consider the sentence ‘Contact: Diane Anderson 076-8268914, info@sandnasbadenscamping.se’ [^2] More examples are in Table 8 in Appendix A) do not provide meaningful content for an NMT system to learn from.
+Improved Taxonomy: Although translation pairs can have overlapping URLs, acronyms, etc, excessive inclusions of such content in a sentence (e.g.
+consider the sentence ‘Contact: Diane Anderson 076-8268914, info@sandnasbadenscamping.se’ [^2] More examples are in Table 11 in Appendix B.) do not provide meaningful content for an NMT system to learn from.
 However, under  Ranathunga et al.
-(2024)’s taxonomy, such sentence pairs would likely be categorized as correct translations.
-Therefore, we define a new noise category CCN to capture such sentence pairs.
-Secondly, we consider the upper limit for short sentences as five words [^3] We consider thresholds 3,4 and 5 and empirically select this threshold as the optimal., based on common observations.
+(2024)’s taxonomy, such sentence pairs would likely be categorised as perfect translation-pairs (CC).
+Therefore, we define a new noise category CCN (high-overlapping non-translatable text) to capture such sentence pairs.
+Secondly, we consider the upper limit for short sentences as five words [^3] We considered thresholds 3,4 and 5 and empirically selected this threshold as it gave the highest NMT gains.
+Results in Table 12 in Appendix C..
 Finally, we improve the definition of WL (wrong language) to consider a threshold in determining whether a sentence pair should be marked as wrong language.
 Table 3 shows these changes.
-The complete list of these noise categories and example parallel sentences is available in Table 9 and Table 10 (respectively) in Appendix A.
-Selection of Heuristics
+The complete list of these noise categories and example parallel sentences are available in Table 9 and Table 10 (respectively) in Appendix B.
+Selection of Heuristics:
 Table 4 shows how the heuristics discussed in Section 2.3 may help in removing different noise categories.
-Note that de-duplication based heuristic cannot be associated with any noise category, as it does not apply to individual sentence pairs.
-In addition to de-duplication strategies discussed in Section 2.3, we introduce an n-gram based de-duplication, meaning that sentences would be removed if they overlap in a consecutive n-gram sequence.
+Note that a deduplication-based heuristic cannot be associated with any in the taxonomy, as it does not apply to individual sentence pairs.
+In addition to the deduplication strategies discussed in Section 2.3, we introduce an n-gram-based deduplication, meaning that sentences would be removed if they overlap in a consecutive n-gram text span.
 Noise Category Short Label Rule-based Heuristic
@@
 Short Sentences CS sLength
-Overlapping Numbers/acronyms/URLs/email CCN LID, sentWRatio, sentCRatio
+High-overlapping non-translatable text CCN LID, sentWRatio, sentCRatio
 Wrong translation X STRatio (With a length difference)
@@
 We conduct a human evaluation to quantify the noise before and after applying the heuristics.
-The annotator selection criteria, resources and training provided, the payment details, etc are described in Appendix F.
-From the top 1000 samples in each of the ranked corpora using LASER3, XLM-R and LaBSE, we randomly select 100 parallel sentences for each language pair.
-We ask each translator to select an annotation category for each sentence pair using the error taxonomy discussed in Section 3.
-Each sentence pair was annotated by two translators to reduce potential bias caused by the individual translators.
-Inter-annotator agreement was calculated using the Pearson correlation coefficient, which resulted in En-Si 0.88, En-Ta 0.68 and Si-Ta 0.49.
+The annotator selection criteria, resources and training provided, the payment details, etc are described in Appendix A.
+For each language-pair, we obtain top 1000 samples in each of the ranked corpora using embeddings obtained from LASER3, XLM-R, and LaBSE, we randomly select 100 parallel sentences for each language pair.
+We ask the translators to annotate each sentence pair using the taxonomy discussed in Section 3.
+Each sentence pair is annotated by three translators to reduce any potential bias inherent in the individual translators.
+For the three annotators, the Fleiss Kappa scores are 0.833 for EnSi, 0.651 for EnTa, and 0.649 for SiTa.
+We note that the results for EnTa and SiTa are very close.
 4 Experiments
@@
 (2020); Ranathunga and de Silva (2022), respectively.
-Languages were selected considering the availability of human evaluators.
+Languages are selected considering the availability of human evaluators.
 We select CCMatrix and CCAligned as the web-mined corpora.
-Both these corpora include parallel data for the language pairs considered in the search.
-More details of these datasets and language pairs are in Appendix B.
-For the NMT experiments, we use the dev and devtest subsets from the Flores+ dataset [^4] https://github.com/openlanguagedata/flores as validation and evaluation sets, respectively.
+Both these corpora include parallel data for the language pairs considered in the research.
+More details of these datasets and language pairs are shown in Appendix D.
+For the NMT experiments, we use the dev and devtest subsets from the Flores-200 Costa-jussà et al.
+(2022) dataset [^4] https://github.com/openlanguagedata/flores as validation and evaluation sets, respectively.
 Dataset statistics are provided in Table 5.
@@
 Table 5: Corpus statistics.
-4.2 Selection of muliPLMs
-We select LASER3, XLM-R, and LaBSE for obtaining embeddings for the sentences to determine the semantic similarity.
-XLM-R, different to others, was trained purely on monolingual data, but has proven to be useful for cross-lingual tasks as well Choi et al.
-(2021); Conneau et al.
-(2020).
-All three models include En, Si, and Ta.
-Details of the multiPLMs are shown in Appendix C.
-4.3 Heuristic-based PDC Experiments
-Each heuristic is applied independently to the source (S), target (T), and both sides (ST).
-In line with the original sentence alignment conducted for CCMatrix and CCAligned, we treat En as the source side.
-For Si-Ta, Si is considered the source because it is more common for a Si sentence to be translated to Ta Farhath et al.
-(2018).
-Finally, for each multiPLM, the retained sentences are ranked in descending order with cosine similarity.
-Deduplication: We consider different granularities of de-duplication.
-ie.
-identical de-duplication (dedup), de-duplicate by removing numbers only (nums) and removing both numbers and punctuations (punctsNums).
-Subsequently, we de-duplicate considering different n-gram spans, i.e.
-4-grams, 5-grams, 6-grams and 7-grams.
-Length-based: We filter short sentences less than five words [^5] We consider sentence length thresholds 3,4 and 5 and empirically decide this value.
-Results in Table 11 in Appendix D.
-While some research has suggested removing extremely long sentences (Minh-Cong et al., 2023; Gala et al., 2023), we found that the percentage of longer sentences is lower and that removing them has a negligible effect.
-Thus, this result is not reported.
-LID-based: We use a public LID model [^6] https://github.com/facebookresearch/fairseq/tree/nllb to predict the language of each sentence.
-The predicted label is then used as a standalone heuristic (LID) and in combination with its associated prediction probability (LIDThresh), with threshold of 0.7777Thresholds below 0.7 reduce NMT results.
-Ratio-based: For STRatio, 0.79-1.39, 0.87-1.62 and 0.85-1.57 were selected as thresholds for En-Si, En-Ta and Si-Ta, respectively.
-These were determined by calculating the mean and the standard deviation obtained for the validation set in a human-crafted En-Si-Ta trilingual dataset (Fernando et al., 2020).
-Following observations of Hangya and Fraser (2018), 0.6 was selected as the threshold for sentWRatio and  sentCRatio.
-4.4 NMT Experiments
-First, a Sentencepiece [^8] https://github.com/google/sentencepiece tokenizer with a vocabulary size of 25000 is trained.
-Then we use the fairseq toolkit  (Ott et al., 2019) to model and train the transformer-based Seq-to-Seq NMT model until convergence.
-Hyperparameters used in the NMT experiments are shown in Table 12 in Appendix E.
-The baseline NMT models are trained on the top 100,000 sentence pairs from the ranked corpus.
-We use ChrF++ Popović (2017) to report NMT results.
-5 Results and Analysis
-We report the results obtained for the NMT models trained in the forward direction.
-The results of the experiments are in Table 6.
-Heuristic-wise best result is summarized in Table 14 in Appendix G.
-As evident from these tables, as well as from Figure 1, NMT results across different multiPLMs show a great disparity in the baseline NMT scores for En-Si and En-Ta language pairs.
-In the following sub-sections, we discuss how the use of heuristics is useful in mitigating this disparity and improving overall NMT results.
-5.1 Impact of Heuristics on NMT Results
 Heuristic Applicable Side Sinhala-Tamil English-Sinhala English-Tamil
@@
 NA would be when the particular experiment is not applicable for that language pair or the dataset.
-5.1.1 Impact of De-duplication based PDC
-We observe that dedup, irrespective of the heuristic-applied side (S/T/ST), outperforms the baseline 89% of the experiments.
-Overall, de-duplication considering both source and target seems to be the most effective - it outperforms the baseline in 94% of the experiments.
-In comparison, dedup target only and source only outperform the baseline in 89% and 83% of the experiments, respectively.
-We apply our newly introduced ngram-based de-duplication dedup+ngram, on top of  dedup.
+4.2 Selection of muliPLMs
+We select LASER3, XLM-R, and LaBSE for obtaining embeddings for the sentences to determine the semantic similarity.
+XLM-R, different to others, was trained purely on monolingual data, but has proven to be useful for cross-lingual tasks as well Choi et al.
+(2021); Conneau et al.
+(2020).
+All three models include En, Si, and Ta.
+Details of the multiPLMs are shown in Appendix E.
+4.3 Heuristic-based PDC Experiments
+Each heuristic is applied independently to the source (S), target (T), and both sides (ST) of the corpora.
+In line with the original sentence alignment conducted for CCMatrix and CCAligned, we treat En as the source side.
+For Si-Ta, Si is considered the source because it is more common for a Si sentence to be translated to Ta Farhath et al.
+(2018).
+Finally, for each multiPLM, the retained sentences are ranked in descending order with cosine similarity.
+Deduplication: We consider different granularities of deduplication.
+i.e.
+identical deduplication (dedup), deduplication after removing numbers only (nums) and removing both numbers and punctuations (punctsNums).
+Subsequently, we deduplicate considering different n-gram spans, i.e.
+4-grams, 5-grams, 6-grams and 7-grams.
+Length-based: We filter short sentences less than five words3.
+While some research has suggested removing extremely long sentences (Minh-Cong et al., 2023; Gala et al., 2023), we find that the percentage of longer sentences is lower and that removing them has a negligible effect.
+Thus, this result is not reported.
+LID-based: We use a public LID model [^5] https://github.com/facebookresearch/fairseq/tree/nllb Costa-jussà et al.
+(2022) to predict the language of each sentence.
+The predicted label is then used as a standalone heuristic (LID) and in combination with its associated prediction probability (LIDThresh), with threshold of 0.7666Thresholds below 0.7 reduce NMT results..
+Ratio-based: For STRatio, 0.79-1.39, 0.87-1.62 and 0.85-1.57 were selected as thresholds for En-Si, En-Ta and Si-Ta, respectively.
+These thresholds are determined by calculating the mean and the standard deviation obtained for the validation set in a human-crafted trilingual dataset (Fernando et al., 2020; Ranathunga et al., 2018).
+Following observations of Hangya and Fraser (2018), 0.6 is selected as the threshold for sentWRatio and  sentCRatio.
+4.4 NMT Experiments
+First, a Sentencepiece [^7] https://github.com/google/sentencepiece tokenizer with a vocabulary size of 25000 is trained.
+Then we use the fairseq toolkit  (Ott et al., 2019) to model and train transformer-based Seq-to-Seq NMT models until convergence.
+Hyperparameters used in the NMT experiments are shown in Table 13 in Appendix F.
+The baseline NMT models are trained on the top 100,000 sentence pairs from the ranked corpus.
+We use ChrF++ Popović (2017) to report NMT results.
+5 Results and Analysis
+We report the results obtained for the NMT models trained in the forward direction.
+The results of the experiments are in Table 6.
+Heuristic-wise best result is summarised in Table 14 in Appendix G.
+As evident from these tables, as well as from Figure 1, NMT results across different multiPLMs show a great disparity in the baseline NMT scores for En-Si and En-Ta language pairs.
+In the following sub-sections, we discuss how the use of heuristics is useful in mitigating this disparity and improving overall NMT results.
+5.1 Impact of Heuristics on NMT Results
+5.1.1 Impact of Deduplication-based PDC
+We observe that deduplication, irrespective of the heuristic-applied side (S/T/ST), outperforms the baseline in 89% of the experiments.
+Overall, dedup considering both source and target seems to be the most effective — it outperforms the baseline in 94% of the experiments.
+In comparison, dedup target-only and source-only outperform the baseline in 89% and 83% of the experiments, respectively.
+Figure 2: Percentage of dedup+ngram experiments exceeding the best result of dedup for each multiPLM
+We apply our newly introduced n-gram-based deduplication dedup+ngram, on top of  dedup.
 We find that for each result column, there is a dedup+ngram result that outperforms the best results obtained with the corresponding dedup result.
-To observe the impact of n value on the NMT result, we plot Figure 2 showing the percentage of ngram experiments exceeding the highest dedup, with respect to the multiPLM (language-wise result is in Figure 4 of the Appendix G).
+To observe the impact of the n value on the NMT result, we plot Figure 2 showing the percentage of n-gram experiments exceeding the highest dedup result, with respect to the multiPLM (language-wise result is in Figure 4 of the Appendix G).
 We observe a consistent pattern - n=5 or 6 perform the best in a majority of cases.
-We believe that 4-gram results in an overly aggressive de-duplication.
-However, the exact ngram depends on the corpus characteristics.
-Figure 2: Percentage of dedup+ngram experiments exceeding the best result of dedup for each multiPLM
+We believe that 4-gram results in an overly aggressive deduplication.
+However, the exact n-gram depends on the corpus characteristics.
 We observe that dedup+punctNums outperforms  dedup+nums and dedup in 78% and 89% of the experiments (respectively), proving that (dedup+punctNums) to be more impactful.
-Finally, we analyse the impact of dedup+puntsNums+ngram999 puntsNums and ngram has been applied on top of dedup..
-Compared to other dedup combinations101010dedup, dedup+ngram, nums, and dedup+punctNums, dedup+punctNums+ngram produce the best result across 67% of the experiments.
+Finally, we analyse the impact of dedup+puntsNums+ngram888puntsNums and ngram has been applied on top of dedup..
+Compared to other dedup combinations999dedup, dedup+ngram, nums, and dedup+punctNums, dedup+punctNums+ngram produce the best result across 67% of the experiments.
 5.1.2 Impact of Length-Based PDC
-sLength surpasses the baseline 89% of the experiments.
+sLength surpasses the baseline in 89% of the experiments.
 Therefore, we can conclude that sLength is favourable as a heuristic.
-When analyzing the side on which the heuristic is applied, we observe that applying it to both ST is the most effective (similar to dedup), followed by T and then S (56%, 28%, and 17% of the experiments, respectively).
-Recall that our manual inspection noticed that XLM-R and LaBSE tend to prioritize shorter sentences.
+When analysing the side on which the heuristic is applied, we observe that applying it to both ST is the most effective (similar to dedup), followed by T and then S (56%, 28%, and 17% of the experiments, respectively).
+Recall that our manual inspection noticed that XLM-R and LaBSE tend to prioritise shorter sentences.
 This observation is affirmed by the sLength results.
@@
 Gains are reported by LASER3 as well, though not very significant.
-The highest gain of +20.61 ChrF++ is reported for XLMR for the CCAligned-EnTa corpus, while a gain of +11.40 is reported by LaBSE for the same corpus.
+The highest gain of +20.61 ChrF++ is reported for XLM-R for the CCAligned-EnTa corpus, while a gain of +11.40 is reported by LaBSE for the same corpus.
 NMT models trained after applying  LID outperform the baseline in 85% of the experiments, while  LIDThresh outperforms LID in 72% of the experiments.
@@
 We observe that the gains for LIDThresh with respect to LID are least with Si-Ta with 50% while for En-Si and En-Ta it is 83% each.
-We assume this is due to a limitation with the LID model, which is not optimized for Si and Ta.
+We assume that this is due to a limitation with the LID model, which is not optimised for Si and Ta.
 5.1.4 Impact of Ratio-based PDC
@@
 5.2 Summary of Heuristic-based PDC
-1.
-Impact of the Individual Heuristics on the NMT Results: The LID-based heuristic emerged as the most impactful in 44% of the experiments.
-De-duplication and sentence-length heuristics were the most effective in 33% and 17% of the experiments, respectively.
+1. Impact of the Individual Heuristics on the NMT Results:
+The LID-based heuristic emerged as the most impactful in 44% of the experiments.
+Deduplication and sentence-length heuristics were the most effective in 33% and 17% of the experiments, respectively.
 The ratio-based heuristic alone did not hold superior results compared to others, making it the least impactful single heuristic.
@@
 For LASER3, XLM-R, and LaBSE, we observe the highest ChrF++ gains of +10.57, +20.61, and +11.40 for CCMatrix-EnTa and CCAligned-EnTa and CCAligned-EnTa corpora, respectively.
-2.
-Impact of Combined Heuristics: The combination of heuristics produced the best score, compared to the best-performing individual heuristic, except for the CCMatrix-SiTa.
-In this, the combination reduced the dataset size by 54% (Table 15 in Appendix G), resulting in only 98k parallel sentences, which could be the reason for the reduced score.
-For CCAligned-SiTa (we exclude CCMatrix due to the dataset size limitation), the gains across multiPLMs are marginal.
+2. Impact of Combined Heuristics:
+The combination of heuristics produced the best score, compared to the best-performing individual heuristics, except for the CCMatrix-SiTa.
+In this, the combination reduced the dataset size by 54% (Table 15 in Appendix G), resulting in only 98k parallel sentences.
+We suspect this reduction in dataset size and the residual noise could result in the reduced score.
+CCMatrix-SiTa gains across multiPLMs are marginal.
 The same pattern holds with LASER3 across En-Ta and En-Si languages, with the exception of CCMatrix-EnTa corpus.
-Since LASER3 is already trained with OPUS [^11] https://opus.nlpl.eu/ data, it is believed that LASER3 is already optimized Moon et al.
+Since LASER3 is already trained with OPUS data Tiedemann and Thottingal (2020) [^10] https://opus.nlpl.eu/, it is believed that LASER3 is already optimised Moon et al.
 (2023) towards ranking the CCMatrix and CCAligned corpora better.
 This is further affirmed with the relatively higher NMT scores of the baseline experiments.
-dedup+punctNums+(n)gram+sLength+LIDThresh combination appears as the best combination in general.
-The specific n-value is dependent on the dataset.
-Only in CCAligned-SiTa with XLM-R, the combination performed best without the LID heuristic.
+For the rest of the ranked corpora, the best NMT scores were reported when the combined heuristic was applied.
+It was noted the combination always had heuristics, dedup+punctNums+(n)gram+sLength+LIDThresh as a common combination for 77% of the experiments.
+The specific n-value was dependent on the dataset.
+In the combination producing the best gains, the ratio-based heuristic was different based on the dataset/language pair.
+In 61% of the cases, the best scores were obtained with sentWRatio, while STRatio yielded the best results in 16% of the experiments.
+The combination performed best without the LID heuristic only in CCAligned-SiTa with XLM-R.
 However, we include LIDThresh into the combination for two reasons.
-(1) Even with LIDThreshold, the score only lags by (-0.19) ChrF++ scores, which is negligible.
-(2) The most influential individual heuristic is LID-based.
-In the combination, producing the best gains, the ratio-based heuristic was different based on the dataset/language pair.
-However, in 80% of cases it had been sentWRatio.
-3.
-Reducing the Disparity Across multiPLMs: In order to calculate the disparity between multiPLMs, we calculated the difference between LASER3 and the other two multiPLMs.
-This is because LASER3 produced the best result for baseline experiments, and its results consistently improved after applying heuristics.
-In Table 7121212Since for the SiTa pair the disparity observed was marginal, we exclude this language-pair from our discussion., the best result of each heuristic category reduces this disparity for all experiments related to CCAlign.
-Disparity has been reduced on average by 95.87% for CCAligned-XLMR/LaBSE and CCMatrix-LaBSE, compared to LASER3.
-The disparity between CCMatrix-XLMR EnSi/EnTa models, compared to LASER3 was reduced by 48% only, meaning that XLM-R ranked corpus contains noise which cannot be mitigated by the heuristics alone.
-However, our hypothesis holds in most cases, and it is safe to say that heuristic-based PDC mitigate the bias brought in by the multiPLM.
+(1) Even with LIDThreshold, the score only lags by (-0.19) ChrF++ scores compared to the best NMT score, which is negligible.
+(2) The most influential individual heuristic is LIDThresh-based for most cases.
+Therefore we recommend the heuristic combination, dedup+punctNums+(n)gram+sLength
++LIDThresh+sentWRatio
+to be applied as the rule-based heuristic combination for the PDC task.
+3. Reducing the Disparity Across multiPLMs:
+We calculate the disparity ($\Delta$) as the difference in NMT scores between LASER3 and the XLM-R or LaBSE scores.
+Equation 1 shows the baseline disparity calculation, where $BL_{\text{{multiPLM}}}$ refers to the baseline NMT score obtained using embeddings from either XLM-R or LABSE during ranking.
+$\Delta_{\textit{Baseline}}=BL_{\text{{LASER3}}}-BL_{\text{{multiPLM}}}$ (1)
+The disparity for each heuristic is calculated with respect to LASER3 and the best NMT score obtained from either XLM-R or LaBSE.
+The results are shown in Table 7.
+Additionally, we calculate the disparity reduction percentage ($\Delta~Reduction(\%)$) as defined in Equation 2.
+Here, $\Delta_{heuristic}$ is the disparity corresponding to the best-performing NMT models after applying the respective heuristic or the optimal combination.
+$\Delta~Reduction(\%)=\frac{\Delta_{\text{baseline}}-\Delta_{\text{heuristic}}}{\Delta_{\text{baseline}}}\times 100\%$ (2)
+In Table 7111111Since the disparity observed was marginal for the SiTa pair, we exclude this language pair from our discussion., we observe that the disparity has been reduced on average by 95.87% for CCAligned-XLM-R/LaBSE and CCMatrix-LaBSE, compared to LASER3.
+The disparity between CCMatrix-XLM-R EnSi/EnTa models, compared to LASER3 was reduced by only 48%, meaning that XLM-R ranked corpus contains noise that cannot be mitigated by the heuristics alone.
+However, our hypothesis holds in most cases, and it is safe to say that heuristic-based PDC mitigates the bias brought in by the multiPLM.
 Heuristic LASER3 vs XLM-R LASER3 vs LaBSE
-Disparity (ChrF++) +/-Disparity wrt Baseline (%) Disparity (ChrF++) +/- Disparity wrt Baseline (%)
+$\Delta$ (ChrF++) $\Delta$ Reduction (%) $\Delta$ (ChrF++) $\Delta$ Reduction (%)
 CCMatrix
@@
 Combined Heuristics 1.73 92.39% 1.45 89.74%
-Table 7: Disparity in ChrF++ points, among the NMT models (XLM-R/LaBSE) with the best scores with respect to LASER3 after applying the individual/combined heuristics.
-The +/- Disparity is this disparity as a percentage of the baseline disparity.
+Table 7: Disparity ($\Delta$) in ChrF++ points, among the NMT models (XLM-R/LaBSE) with the best scores with respect to LASER3 after applying the individual/combined heuristics.
+The $\Delta$ Reduction (%) is this disparity as a percentage of the baseline disparity.
 5.3 Human Evaluation Results
 As shown in Table 2, heuristic-based PDC had reduced the noise in the top-ranked samples consistently, irrespective of the language pair and the considered multiPLM.
-Some gains are quite significant - For example, the amount of correct pairs (C) improvement for CCMatrix-EnSi (LaBSE), CCAligned-EnTa (LaBSE) and CCAligned-EnSi (XLM-R) were all 82%.
-Before applying the heuristics, the average Error category (E) across language pairs and datasets is 13.50% for LASER3, 36.33% for XLM-R and 72.53% for LaBSE.
-CS, CCN and UN noise categories were noted to be contributing towards this noise percentage.
+Some gains are quite significant — for example, the amount of correct pairs (C) improvement for CCMatrix-EnSi (LaBSE), CCMatrix-EnTa (LaBSE), and CCAligned-EnSi (LaBSE) were 82%, 82%, and 74% respectively.
+CS, CCN and UN noise categories are noted to be contributing towards this noise percentage.
 However, after heuristic filtration, the error drops drastically.
-For LASER, XLM-R and LaBSE on average, it is 2.17%, 2.50% and 1.0%, respectively.
-However, the untranslated (UN) and overlapping text (CCN) were still there.
-To eliminate these types of noise, we would need to employ an alignment model similar to Steingr´ımsson et al.
+On average, for LASER, XLM-R, and LaBSE, the final error percentages are 2.17%, 2.50% and 1.0%, respectively.
+The evaluation further reveals that the residual noise are from untranslated (UN) and overlapping text (CCN) classes.
+These findings reveal that rule-based heuristics, are not effective in eliminating those types of noise.
+As a result, we would need to employ an alignment model similar to Steingrímsson et al.
 (2023) or Minh-Cong et al.
-(2023).
-In conclusion, the human evaluation results also indicate that the heuristic-based PDC approach is beneficial for multiPLM ranking in two key ways.
+(2023) to remove such residual noise from the corpus.
+In conclusion, the human evaluation results indicate that the heuristic-based PDC approach is beneficial for parallel sentence ranking in two key ways.
 First, it produces the top-ranked sentence pairs from multiPLM to be qualitatively comparable.
@@
 We further investigate the effect of the training dataset size on the final NMT performance.
-We sample 50k, 100k, 150k, and 200k from each ranked corpus, and train NMT models.
+We sampled the top 50k, 100k, 150k, and 200k from each of the ranked corpora after applying the heuristic-combination, and train NMT models.
 Results are in Figure 5 of the Appendix H.
-For Si-Ta, moderate but consistent improvements were observed from 50K to 100K samples for both CCMatrix and CCAligned ranked corpora, irrespective of the multiPLMs.
-However, due to the lack of sufficient corpus size, the NMT scores for 150K and 200K were not observable.
-For En-Ta, NMT scores declined after 100K, while for En-Si, an improvement was observed until 200K, except for the CCMatrix XLM-R ranked sample.
+For Si-Ta, moderate but consistent improvements are observed from 50K to 100K samples for both CCMatrix and CCAligned corpora, irrespective of the multiPLMs.
+However, after heuristic-based filtering, the resulting Si-Ta parallel corpus contained only about 100K sentence pairs (Table 15 in Appendix G).
+Therefore, NMT experiments with 150K and 200K sentence pairs could not be conducted due to insufficient data.
+For En-Ta, NMT scores declined after 100K, while for En-Si, an improvement was observed until 200K, except for the CCMatrix ranked using XLM-R embeddings.
 Hence, we suspect that noise may be more prominent beyond 100K for En-Ta.
 In conclusion, increasing corpus size can enhance NMT performance, but only when the underlying data is sufficiently clean and well-aligned.
-7 Discussion
-We show that applying simple rule-based heuristics to remove noisy sentences before using a multiPLM is better than simply ranking the web-mined corpora based on embeddings derived from multiPLMs as suggested by Ranathunga et al.
-(2024).
-Based on our extensive experiments, dedup+punctNums+(n)gram+sLength+LIDThresh
-+sentWRatio is our recommended heuristic combination, based on the justifications provided earlier.
-Although $n$ can be corpus dependent, we recommend starting with n=5 or 6.
-In contrast to the recent PDC work Steingr´ımsson (2023); Minh-Cong et al.
-(2023) that employs several deep learning model-based rigorous filtration pipelines, our technique is much simpler.
 7 Conclusion
@@
 Therefore, for anyone planning to use web-mined corpora, our recommendation is to first filter out noisy sentences using heuristics and then to do ranking on the embeddings derived from the multiPLM.
+In contrast to the recent PDC work Steingrimsson (2023); Minh-Cong et al.
+(2023) that employs several deep learning model-based rigorous filtration pipelines, our technique is much simpler.
 Human evaluation indicated that even after applying heuristics, some noise remains.
-In future, we plan to apply techniques such as classification-based approaches to remove such noise, or to apply translation post-editing to improve sentence quality.
+In future, we plan to apply techniques such as classification-based approaches to remove such noise, or to apply translation post-editing to improve the sentence quality.
 8 Limitations and Ethical Concerns
@@
 Therefore, such models would need to be optimised or better LID models would need to be used, such that they are effective in predicting the language label correctly.
-Due to the lack of funding, we were able to employ only two annotators for the task.
-However, the results would have been more convincing if the annotation could have been done by three annotators.
 Furthermore, due to the lack of human annotators, we had to limit this analysis to only three language pairs.
@@
 We have offered the annotators the standard rate and have settled the payments in full.
-The only details we disclose are in the Appendix F.
+The only details we disclose are in Appendix A.
 Acknowledgments
-Will be provided upon acceptance.
-Appendix A Improved Noise Taxonomy
-The definitions of all the noise categories in our improved noise taxonomy are given in Table 9. Then, in Table 10, we show examples of parallel sentences which fall under these noise categories.
-Table 8: Example parallel sentences which will be separately identified under the new noise category CCN
-Error Code: Description
+This research was funded by the Google Award for Inclusion Research (AIR) 2022 received by Surangika Ranathunga and Nisansa de Silva.
+We would also like to thank and acknowledge the National Languages Processing Centre (NLPC), at the University of Moratuwa for providing the GPUs to execute the experiments related to the research.
+Appendix A Human Evaluation
+Selection of the Annotators: We select annotators for this task who have translation or machine translation-related experience for a minimum of 2 years and are fluent in both languages of the specific language pair assigned to them.
+Table 8 shows the years of experience and the qualifications of those annotators who conducted this task.
+Annotator Experience Qualification
+Annotator 01 22 Diploma in Translation And Interpretation
+Annotator 02 9 BSc.(Hons) in Information Technology
+Annotator 03 5 MBBS
+Annotator 04 4 BA (Hons) in Translation
+Annotator 05 3 BSc (Hons) Engineering sp. in Computer Science and Engineering
+Annotator 06 2.5 Diploma in Translation And Interpretation
+Annotator 07 2.5 BSc Eng (Hons) Electrical & Electronics Engineering
+Annotator 08 2.5 BSc (Hons) Engineering sp. in Electrical Engineering
+Annotator 09 2 Bachelor of Industrial Information Technology
+Table 8: Annotator details with the years of experience and their qualifications.
+Resources Provided and Training: All the annotators have had prior experience with a similar task Ranathunga et al.
+(2024).
+However, for this annotation work, we provide them with the definitions of the noise categories (Table 9) along with example sentence pairs (Table 10) and the guideline in terms of a flowchart (Figure 3).
+First, we asked them to do a sample of 30 sentences as a training on the task, and review it.
+Then, the 1200 sentence pairs to be annotated were shared with each annotator via Google Sheets to be completed.
+Parallel Corpus Categorization Code: Description
 CC: Perfect Translation-pair
@@
 Less than 5 words. Translation-wise, correct, but only a short phrase or a few words.
-CCN: Overlapping Content
+CCN: High-overlapping non-translatable text
 Perfect or near-perfect translation pair, but with overlapping content like numbers, acronyms, or URLs. Sentences longer than 5 words with high overlap.
@@
 Either the source or the target (or both) is not in the expected language. Up to 30% of acceptable content may be tolerated.
-Table 9:  Ranathunga et al. (2024)’s error taxonomy with the CCN category that has been newly added by us
-Table 10: Example parallel sentences from the En-Si, En-Ta and Si-Ta, identified during human evaluation.
-Appendix B Selection of Languages and Datasets
-This section contains details on the selected languages and the web-mined corpora considered under the study.
-Sinhala is an Indo-Aryan language spoken primarily in Sri Lanka by the Sinhalese majority.
-It exhibits complex morphological structures, including rich inflectional and derivational processes, but is classified as a low-resource language due to the scarcity of linguistic resources and tools.
-Tamil, a Dravidian language with a rich literary history, is spoken by Tamil communities in Sri Lanka, India, and the global diaspora.
-Unlike Sinhala, Tamil benefits from a relatively larger digital presence, but it still faces challenges in NLP applications due to morphological complexity, agglutinative grammar, and resource limitations in certain domains.
-We obtain the publicly released CCMatrix and CCAligned datasets from the OPUS website [^13] https://opus.nlpl.eu/.
-Both these datasets support the language pairs, En-Si, En-Ta and Si-Ta, which we consider in this research.
-CCMatrix Schwenk et al.
-(2021) is a web-mined parallel corpus extracted using LASER2-based sentence embeddings to align bitext.
-While it provides large-scale data, it is highly noisy due to the global mining approach to determine alignments, resulting in misaligned or low-quality translations.
-CCAligned El-Kishky et al.
-(2020) extracts bitext from Common Crawl [^14] https://commoncrawl.org/ using document-level and sentence-level alignment based on multilingual embeddings.
-Though it improves alignment quality over global bitext-mined corpora, it still contains significant noise, requiring careful filtering for reliable use.
-Appendix C Selection of multiPLMs
-We include the details on the three multiPLMs considered in this study.
-LASER3 Heffernan et al.
-(2022) (L=12, H=1024, A=4, 250M) [^15] No of Layers, Hidden Layer Dimensions and No of Attention Heads are defined by L, H and A respectively.
-is a multiPLM favourable for bitext mining and cross-lingual tasks.
-It improves over previous LASER2 versions by supporting more languages and enhancing alignment quality, but still faces challenges in low-resource settings.
-XLM-R Conneau et al.
-(2020) (L=12, H=768, A=6, 278M parameters) is a transformer-based multiPLM trained on massive amounts of text using masked language modelling.
-It achieves strong cross-lingual performance but struggles with low-resource languages due to limited training data.
-LaBSE Feng et al.
-(2022) (L=12, H=768, A=12, 471M) is a BERT-based model optimized for multilingual sentence embeddings and bitext retrieval.
-It provides high-quality cross-lingual representations and is favourable for cross-lingual tasks.
-Appendix D Sentence Length Threshold
-To determine the optimal threshold for sentence length filtering, we filter sentences less than 3, 4, and 5 words on the source (S) side, target (T) side, and both sides.
-Similar to other PDC experiments, we then rank the parallel sentence pairs by computing embeddings using three different multiPLMs, and train NMT models using the top 100k ranked data.
-We conduct these experiments using the CCAligned corpus, for En-Si and En-Ta language-pairs.
+Table 9: Improved  Ranathunga et al. (2024)’s taxonomy to categorize parallel sentences to identify biases in multiPLMs.
+Table 10: Example parallel sentences from the En-Si, En-Ta and Si-Ta, identified during human evaluation. The translation error in the language pair is highlighted in pink.
+Figure 3: Shows the annotation guideline document in terms of a flow chart. This shows the priority of the noise category to be selected prior to declaring the annotation class.
+Compensation: They were paid the standard rate in Sri Lanka for each sentence pair they annotated or were offered co-authorship of this paper.
+Table 11: Example parallel sentences which will be separately identified under the new noise category CCN
+Appendix B Improved Noise Taxonomy
+In this research, we extend the taxonomy of  Ranathunga et al.
+(2024) by introducing the noise category CCN (high-overlapping non-translatable text) to capture the type of noise that multiPLMs are biased to rank highly.
+Table 11 shows examples for this noise type.
+The definitions of all the categories in our improved taxonomy are given in Table 9.
+Then, in Table 10, we show examples of parallel sentences which fall under each of these categories.
+Appendix C Sentence Length Threshold
+To determine the optimal threshold for sentence length filtering, we filter sentences with fewer than 3, 4, and 5 words on the source (S) side, target (T) side, and both sides.
+Similar to other PDC experiments, we then rank the parallel sentence pairs by computing embeddings using XLM-R, LABSE and LASER.
+Finally, we train NMT models using the top 100k-ranked data.
+We conduct these experiments using the CCAligned corpus for En-Si and En-Ta language pairs.
 The ChrF++ results are shown in Table 12.
-We observe consistently sentence length threshold of 5 yields the best results.
+We observe sentence length threshold of 5 yields the best results consistently.
 Therefore, we select this threshold in all subsequent experiments.
@@
 Table 12: NMT results in ChrF++ for different sentence length thresholds.
-Appendix E NMT Experiments
-The experiments were conducted on a NVIDIA Quadro RTX6000 GPU with 24GB VRAM.
+Appendix D Selection of Languages and Datasets
+This section contains details on the selected languages and the web-mined corpora considered under the study.
+Sinhala
+is an Indo-Aryan language spoken primarily in Sri Lanka by the Sinhalese majority.
+It exhibits complex morphological structures, including rich inflectional and derivational processes, but is classified as a low-resource language due to the scarcity of linguistic resources and tools de Silva (2025); Ranathunga and de Silva (2022).
+Tamil,
+a Dravidian language with a rich literary history, is spoken by Tamil communities in Sri Lanka, India, and the global diaspora.
+Unlike Sinhala, Tamil benefits from a relatively larger digital presence, but it still faces challenges in NLP applications due to morphological complexity, agglutinative grammar, and resource limitations in certain domains Wijeratne et al.
+(2019).
+We obtain the publicly released CCMatrix and CCAligned datasets from the OPUS collection Tiedemann and Thottingal (2020) [^12] https://opus.nlpl.eu/.
+Both these datasets support the language pairs, En-Si, En-Ta and Si-Ta, which we consider in this research.
+CCMatrix Schwenk et al.
+(2021) is a web-mined parallel corpus extracted using LASER2-based sentence embeddings to align bitext.
+While it provides large-scale data, it is highly noisy due to the global mining approach to determine alignments, resulting in misaligned or low-quality translations.
+CCAligned El-Kishky et al.
+(2020) extracts bitext from Common Crawl [^13] https://commoncrawl.org/ using document-level and sentence-level alignment based on multilingual embeddings.
+Though it improves alignment quality over global bitext-mined corpora, it still contains significant noise, requiring careful filtering for reliable use.
+Appendix E Selection of multiPLMs
+We include the details on the three multiPLMs considered in this study.
+LASER3 Heffernan et al.
+(2022) (L=12, H=1024, A=4, P=250M) [^14] No.
+of Layers, Hidden Layer Dimensions, No of Attention Heads, and Total number of parameters are defined by L, H, A, and P respectively.
+is a multiPLM favourable for bitext mining and cross-lingual tasks.
+It improves over previous LASER2 versions by supporting more languages and enhancing alignment quality, but still faces challenges in low-resource settings.
+XLM-R Conneau et al.
+(2020) (L=12, H=768, A=6, P=278M) is a transformer-based multiPLM trained on massive amounts of text using masked language modelling.
+It achieves strong cross-lingual performance but struggles with low-resource languages due to limited training data.
+LaBSE Feng et al.
+(2022) (L=12, H=768, A=12, P=471M) is a BERT-based model optimised for multilingual sentence embeddings and bitext retrieval.
+It provides high-quality cross-lingual representations and is favourable for cross-lingual tasks.
+Appendix F NMT Experiments
+The experiments are conducted on a NVIDIA Quadro RTX6000 GPU with 24GB VRAM.
 The hyperparameters used during training, along with the training parameters, are shown in Table 13.
@@
 ChrF++ was chosen over the conventional multi-BLEU Papineni et al.
-(2002) and sacreBLEU (Post, 2018) because character-level evaluation is more suitable for the considered Sinhala and Tamil languages, which are morphologically rich in nature.
+(2002) and sacreBLEU (Post, 2018) because character-level evaluation is more suitable for the considered languages, Sinhala and Tamil, which are morphologically rich in nature.
 Hyperparameter Argument value
@@
 Table 13: Training parameters for NMT experiments.
-Appendix F Human Evaluation
-Selection of the Annotators: We select annotators for this task who have translation or machine translation-related experience for a minimum of 2 years and are fluent in both languages of the specific language pair assigned to them.
-Table 13 shows the years of experience and the qualifications of those annotators who conducted this task.
-Annotator Experience Qualification
-Annotator 01 22 Diploma in Translation And Interpretation
-Annotator 02 9 BSc.(Hons) in Information Technology
-Annotator 03 5 MBBS
-Annotator 04 4 BA (Hons) in Translation
-Annotator 05 3 BSc (Hons) Engineering sp. in Computer Science and Engineering
-Annotator 06 2.5 Diploma in Translation And Interpretation
-Annotator 07 2.5 BSc Eng (Hons) Electrical & Electronics Engineering
-Annotator 08 2.5 BSc (Hons) Engineering sp. in Electrical Engineering
-Annotator 09 2 Bachelor of Industrial Information Technology
-Table 13: Annotator details with the years of experience and their qualifications.
-Resources Provided and Training: All the annotators have had prior experience with a similar task.
-However, for this annotation work, we provide them with the definitions of the noise categories (Table 9) along with example sentence pairs (Table 10) and the guideline in terms of a flowchart (Figure 3).
-First, we asked them to do a sample of 30 sentences as a training on the task, and review it.
-Then, the 1200 sentence pairs to be annotated were shared with each annotator via Google Sheets to be completed.
-Compensation: They were paid for each sentence pair they annotated, which was the standard rate for such work.
-Appendix G Results Analysis
-We show the individual and the combined heuristic which produced the best NMT gains with respect to each multiPLM, dataset and the language pair in the Table 14.
-The Figure 4 shows the percentage of dedup+ngram experiments exceeding the highest dedup, with respect to the language-pair.
-Finally, we show the final dataset sizes after applying the heuristic along with the percentage reduction in Table 15.
 Heustic/s CCMatrix Heustic/s CCAligned
@@
 Here DD+PN is Deduplication+punctNums, SL is sLength and LT is LIDThresh.
+Figure 4: Percentage of dedup+ngram experiments exceeding the best result of dedup with respect to the Language-pair.
+(a) En-Si
+(b) En-Ta
+(c) Si-Ta
+Figure 5: ChrF++ scores of NMT systems trained by varying the training dataset size.
+Appendix G Results Analysis
+We show the individual and the combined heuristic which produced the best NMT gains with respect to each multiPLM, dataset and the language pair in the Table 14.
+The Figure 4 shows the percentage of dedup+n-gram experiments exceeding the highest dedup, with respect to the language-pair.
+Finally, we show the final dataset sizes after applying the heuristic along with the percentage reduction in Table 15.
 Heuristic Applicable Side Sinhala - Tamil English - Sinhala English - Tamil
@@
 NA corresponds to the experiments that are not applicable for the language-pair.
-Figure 3: Shows the annotation guideline document in terms of a flow chart. This shows the priority of the noise category to be selected prior to declaring the annotation class.
-Figure 4: Percentage of dedup+ngram experiments exceeding the best result of dedup with respect to the Language-pair.
 Appendix H NMT Performance on Training Dataset
 Figure 5 shows the variation in NMT results by varying the dataset size with 50k, 100, 150k and 200k for each language-pair.
-(a) En-Si
-(b) En-Ta
-(c) Si-Ta
-Figure 5: ChrF++ scores of NMT systems trained by varying the training dataset size.